### v3

# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time effects in both the intercept and slope. To address the incidental parameter problem inherent in this class of models, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting maximum likelihood estimators are automatically bias-corrected. Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-square distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
Keywords:
incidental parameter problem; bias correction; fixed effects; maximum likelihood
JEL Classification:
C23
1 Introduction
Panel data models with two-way fixed effects are widely used in empirical research to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007). When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
To illustrate, consider the classic problem of estimating the impact of children on a mother’s labor force participation. Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins 1988; Angrist and Evans 1998; Kleven et al. 2019). First, mothers who choose to have more children might be inherently less inclined to work. Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or access to childcare. In a panel dataset tracking households over time, we can control for the first type of level heterogeneity by including fixed effects in the intercept. However, to address the second type of response heterogeneity, we must include fixed effects in the slope. Given that labor force participation is a binary outcome and typically exhibits serial correlation, this example underscores the need for a dynamic nonlinear panel data model with fixed effects in both the intercept and slope.
In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models. To address the incidental parameter problem inherent in DN-HP models, we propose a likelihood-based analytical bias correction approach. This method ensures the validity of the maximum likelihood estimator and likelihood-based inference for the structural parameters of DN-HP models. As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing maximum likelihood estimators from our bias-corrected likelihoods and the original likelihoods. The results reveal that the maximum likelihood estimators obtained from the bias-corrected likelihoods significantly reduce estimation bias while maintaining comparable mean square errors.
Figure 1: Comparison of Uncorrected and Corrected Maximum Likelihood Estimators
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it},$ and the true values $\rho_{0}=\beta_{0}=0.5$. ($\widehat{\beta},\widehat{\rho})$ are the uncorrected estimators and $(\widehat{\beta}_{L},\widehat{\rho}_{L})$ are the corrected estimators. $1000$ replications. For details on the data-generating process, please refer to Section 4.
1.1 Literature Review
The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity. Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept. To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework. For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator. Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods. Woutersen (2002) and Li et al. (2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood. Most of these methods are analytical, estimating the bias using explicit formulas. Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of $O(1/N)$. Moon and Weidner (2017) propose a parameter-based analytical correction for dynamic linear models with two-way interactive effects. For nonlinear models, Fernández-Val and Weidner (2016) develop parameter-based bias correction techniques with additive fixed effects, while Chen et al. (2021) focus on interactive fixed effects in static models. Alternatively, Jochmans and Otsu (2019) propose a likelihood-based analytical bias correction for static models, allowing fixed effects to enter the model in an arbitrary but known form. Other model-specific solutions include Bai (2009) and Charbonneau (2017), among others. For an extensive review of the incidental parameter problem for models with two-way fixed effects in the intercept, see Fernández-Val and Weidner (2018).
To our knowledge, our paper represents one of the first efforts to extend intercept-only two-way fixed effects models into multi-dimensional two-way fixed effects models, accounting for both level and response heterogeneities. Two prior studies in this line of research are Keane and Neal (2020) and Lu and Su (2023). Both examine linear heterogeneous parameter models with additive effects in both the intercept and slope. Keane and Neal (2020) propose an iterative procedure, which they refer to as the “mean observation OLS,” to eliminate the bias, while Lu and Su (2023) rely on a parameter-based Jackknife approach and provide a uniform inference theory for structural parameters. Our study significantly expands the model framework and introduces a likelihood-based analytical bias correction technique to ensure valid estimation and inference. Neither our DN-HP models nor our bias correction procedure have been discussed in the current literature.
1.2 This Paper
We assume that the multi-dimensional fixed effects enter the log-likelihood additively. Identification restrictions are imposed on the fixed effects, whose dimension grows with the sample size. Our corrected likelihood is derived using two Taylor expansions: the expansion of the score function and the expansion of the original likelihood, both with respect to the fixed effects. Under appropriate dominance conditions, we explicitly derive the two additive bias terms in the original likelihoods of DN-HP models. These bias terms are expressed as quadratic forms involving the score and Hessian of the fixed effects. We estimate the derived bias terms under weak serial dependence (conditional strong mixing) and conditional cross-sectional independence.
We rigorously show that the maximum likelihood estimator (MLE) based on the bias-corrected likelihood of the DN-HP model retains a limiting normal distribution with zero mean (automatically bias-corrected), consistent with the limiting distribution of the classical MLE in the absence of incidental parameters. Beyond the bias-corrected MLE, this likelihood-based bias correction approach also facilitates hypothesis testing using the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics. We demonstrate that these three test statistics, derived from the bias-corrected likelihood, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife. In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
The remainder of the paper is organized as follows. Section 2 describes our settings, the identification restriction, the incidental parameter problem under DN-HP, and gives an overview of our bias correction procedures. Section 3 provides the bias expansion of the likelihood and the formula for our corrected likelihood. The statistical properties of the corrected likelihood. We rigorously show that the estimator of the structural parameter, and the LR, LM, and Wald statistics all follow their standard asymptotic distribution. Section 4 presents simulation studies to demonstrate our method’s performance. Here we show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors. In addition, we also show that our procedure restores the test size of the LR test to its nominal level. We also compare our procedure with the Jackknife and find that our procedure outperforms the Jackknife. Section 5 applies our method to study single mothers’ labor force participation. Proofs, technical details, and additional results are provided in the appendix.
Notations. In what follows, we denote $\iota_{n}$ to be an $n\times 1$ vector of ones and for $\mathbb{I}_{n}$ to be an $n\times n$ identity matrix. Let $\otimes$ be the kronecker product and $\mathds{1}\left\{\cdot\right\}$ the indicator function. In addition, for each square matrix $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\left\{A_{1},\ldots,A_{n}\right\}$ represent a block diagonal matrix, where each matrix $A_{i}$ is the $i$-th diagonal element.
2 Bias in Heterogeneous Parameter Model
2.1 DN-HP Model
Consider the responses $Y_{it},$ which follow the conditional density $f\left(Y_{it};\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}\right),$ given covariates $X_{it}$ (omitted for simplicity). Here, $i=1,\ldots,N$ represents individuals, and $t=1,\ldots,T$ represents time periods. The density is known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0},$ and the covariates $X_{it}$ may include lagged $Y_{it}$. Specifically, $\theta_{0}$ is the $K$-dimensional structural parameter of interest, where $K$ is fixed and known. The nuisance heterogeneous effects $\alpha_{i}^{0}=\left(\alpha_{1,i}^{0},\ldots,\alpha_{K,i}^{0}\right)^{\prime}$ and $\gamma_{t}^{0}=\left(\gamma_{1,t}^{0},\ldots,\gamma_{K,t}^{0}\right)^{\prime}$ may be correlated with $X_{it}$. Throughout the paper, we treat $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ as fixed effects (unknown parameters) by conditioning on them in the density.
Denote $\phi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime}\right)^{\prime}$ and define the quasi-likelihood function as
$l(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}\left(\theta,\phi\right),\qquad l_{i,t}(\theta,\phi):=\log f\left(Y_{it};\theta+\alpha_{i}+\gamma_{t}\right).$
Note that if the data are independent across both $i$ and $t,$ $l(\theta,\phi)$ represents the exact likelihood. However, when observations for each unit $i$ exhibit serial correlation, $l(\theta,\phi)$ is misspecified even though the likelihood for each observation $Y_{it}\mid X_{it},$ $l_{i,t}(\theta,\phi),$ is correctly specified. Our asymptotic theory assumes conditional cross-sectional independence and weak time-serial dependence (Assumption 1). For simplicity, we refer to $l(\theta,\phi)$ as the likelihood function hereafter. Here is an example that illustrates how our model captures greater heterogeneity than conventional two-way effects panel data models.
Example 1 (heterogeneous parameter model)
Consider a probit model
$Y_{it}=\mathds{1}\left\{\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)+\varepsilon_{it}>0\right\}$
where the linear index $\pi_{it}(\cdot)$ varies across both individuals and time periods. The error term $\varepsilon_{it}$ follows a standard normal distribution denoted by $\Phi$. Without of loss of generality, let $X_{it}\in\mathbb{R},$ and the parameter vectors $\theta=\left(\theta_{1},\theta_{2}\right)^{\prime},$ $\alpha_{i}=\left(\alpha_{1,i},\alpha_{2,i}\right)^{\prime},$ and $\gamma_{t}=\left(\gamma_{1,t},\gamma_{2,t}\right)^{\prime}.$ We accommodate the linear index of the form:
$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\left(\theta_{2}+\alpha_{2,i}+\gamma_{2,t}\right).$ (2.1)
In the special case where $\alpha_{i}=\alpha_{1,i}$ and $\gamma_{t}=\gamma_{1,t},$ (2.1) simplifies to
$\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(\theta_{1}+\alpha_{1,i}+\gamma_{1,t}\right)+X_{it}\theta_{2},$
which corresponds to the traditional two-way fixed effects model. Assume that $\varepsilon_{it}$ is independent of $X_{it}.$ The likelihood for individual $i$ at time $t$ is given by
$l_{i,t}\left(\theta,\phi\right):=Y_{it}\log\Phi\left(\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right)+\left(1-Y_{it}\right)\log\Phi\left(-\pi_{it}\left(\theta,\alpha_{i},\gamma_{t}\right)\right).$
Due to the additive form of the parameters, the structural parameters, $\theta,$ is not identified in the likelihood. To resolve this, we reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t},$ thereby imposing the constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0.$ Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}.$ The reparameterized likelihood is then $l\left(\theta,D^{\prime}\psi\right)=:l^{*}\left(\theta,\psi\right)$ for $D:=\operatorname*{diag}\{D_{1},D_{2}\}$ where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}.$
2.2 Incidental Parameter Problem
In this section, we illustrate the incidental parameter problem (IPP) that arises in our heterogeneous parameter model. For a given $\theta,$ define
$\psi(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}\mathbb{E}[l^{*}(\theta,\psi)],\quad\widehat{\psi}(\theta):=\arg\max_{\psi\in\Lambda_{\theta}}l^{*}(\theta,\psi)$ (2.2)
where $\Lambda_{\theta}$ is a compact subset of $\mathbb{R}^{(N+T-2)K}$ for fixed $N$ and $T$ (Assumption 1) and $\mathbb{E}$ denotes expectations with respect to the true conditional density. Specifically, $\mathbb{E}$ represents the conditional expectation given the true values of fixed effects and the covariates. For notational simplicity, we use $\mathbb{E}$ when no confusion arises. Let the score and Hessian be
$s\left(\theta,\psi\right):=\partial_{\psi}l^{*}\left(\theta,\psi\right),\quad H\left(\theta,\psi\right):=\partial_{\psi\psi^{\prime}}l^{*}\left(\theta,\psi\right),$
where $\partial_{\psi}$ and $\partial_{\psi\psi^{\prime}}$ represent the first and second partial derivatives with respect to $\psi$. By (2.2), we know that $s(\theta,\widehat{\psi}\left(\theta\right))=0.$ Given $\theta,$ consider a Taylor expansion of $s(\theta,\widehat{\psi}\left(\theta\right))$ around $\psi(\theta)$:
$0=s\left(\theta,\psi\left(\theta\right)\right)+H\left(\theta,\psi(\theta)\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{s}(\theta,a),$ (2.3)
where
$r^{s}\left(\theta,a\right)$ $:=\big{(}r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\big{)}^{\prime},$
$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\big{(}\widehat{\psi}\left(\theta\right)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)},$
$G_{\psi_{j}}(\theta,a_{j})$ $:=\partial_{\psi_{j}}H(\theta,a_{j}\widehat{\psi}(\theta)+(1-a_{j})\psi\left(\theta\right))$
with $a=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1),$ and $\psi_{j}$ denotes the $j$th component of $\psi.$ Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j},$ evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right).$ For notational simplicity, we write
$s(\theta):=s\left(\theta,\psi\left(\theta\right)\right)\quad\text{and}\quad H\left(\theta\right):=H\left(\theta,\psi(\theta)\right).$ (2.4)
From (2.3), by Assumption 1, we obtain that
$\widehat{\psi}(\theta)-\psi(\theta)$ $=[-{H}(\theta)]^{-1}\big{(}s(\theta)+r^{s}\left(\theta,a\right)\big{)}.$ (2.5)
Next, define the profile likelihoods as $\widehat{l}(\theta):=l^{*}(\theta,\psi)\mid_{\psi=\widehat{\psi}(\theta)}$ and $l(\theta):=l^{*}(\theta,\psi)\mid_{\psi={\psi}(\theta)}$. A Taylor expansion of $\widehat{l}(\theta)$ around $\psi(\theta)$ gives
$\widehat{l}(\theta)=l(\theta)+s^{\prime}\left(\theta\right)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}{H}(\theta)\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}+r^{l}\left(\theta,a_{0}\right)$ (2.6)
where $r^{l}\left(\theta,a_{0}\right):=1/3\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{(N+T-2)K})$ for some $a_{0}\in(0,1).$ Combining (2.5) and (2.6), we have
$\widehat{l}(\theta)=l(\theta)-\frac{1}{2}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)+r(\theta)$ (2.7)
where $r(\theta):=1/2(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)+r^{l}\left(\theta,a_{0}\right).$ Here we suppress $a$ and $a_{0}$ in $r(\theta)$ as they are constants. Under mild regularity conditions, we show that
$\sqrt{NT}\mathbb{E}\big{\{}s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\big{\}}\nrightarrow 0$
and that $r(\theta)$ is a remainder term that is negligible for inference (see Section 3). This implies that the presence of $s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)$ in (2.7), due to the estimation of $\psi(\theta),$ contributes the MLE of $\theta_{0}$ based on $\widehat{l}(\theta).$ Consequently, the structural estimators and $\widehat{l}(\theta)$ based test statistics no longer follow the standard asymptotic distributions as in classical maximum likelihood theory. This is referred to as the IPP in this context. We use the following example to illustrate IPP.
Example 2 (heterogeneous means)
Assume that $Y_{it}$ follows a normal distribution with mean $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and unit variance, independent across both $i$ and $t.$ $\theta_{0}$ is the structural parameter of interest, while $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ are fixed effects. The likelihood function is
$l\left(\theta,\phi\right)=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}-\frac{1}{2}\left(Y_{it}-\theta-\alpha_{i}-\gamma_{t}\right)^{2}-\frac{1}{2}\log(2\pi).$
Define $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}.$ Under the identification condition, the reparameterized likelihood $l^{\ast}(\theta,\psi)=l(\theta,D^{\prime}\psi).$ Given $\theta,$ maximizing $l^{\ast}(\theta,\psi)$ with respect to $\psi$ yields
$\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0}$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad i=1,\ldots,N-1,$ (2.8)
$\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\quad t=1,\ldots,T-1.$ (2.9)
Let $\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta)\big{)}^{\prime}=:\widehat{\psi}(\theta)$ and $\big{(}{\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0}\big{)}^{\prime}=:\psi_{0}.$ Define $\widehat{\alpha}_{N}(\theta):=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}(\theta)$ and $\widehat{\gamma}_{T}(\theta):=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}(\theta).$ Note that (2.8) and (2.9) also hold for $i=N$ and $t=T$ respectively. Denote $\phi_{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\alpha}_{N}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0},{\gamma}_{T}^{0})^{\prime}$ and $\widehat{\phi}(\theta):=\big{(}\widehat{\alpha}_{1}(\theta),\ldots,\widehat{\alpha}_{N-1}(\theta),\widehat{\alpha}_{N}(\theta),\widehat{\gamma}_{1}(\theta),\ldots,\widehat{\gamma}_{T-1}(\theta),\widehat{\gamma}_{T}(\theta)\big{)}^{\prime}.$ Note that
$\widehat{l}(\theta)-l(\theta)$ $=$ $l(\theta,\widehat{\phi}(\theta))-l(\theta,\phi_{0})$
$=$ $\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})\big{]}(\widehat{\alpha}_{i}(\theta)-\alpha_{i}^{0})$
$+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\big{[}Y_{it}(\theta)-\frac{1}{2}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})\big{]}(\widehat{\gamma}_{t}(\theta)-\gamma_{t}^{0})$
where $Y_{it}(\theta):=Y_{it}-\theta-\alpha_{i}^{0}-\gamma_{t}^{0}.$ Since $Y_{it}$ is independent across $i$ and $t,$ by substituting (2.8) and (2.9), we obtain
$\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}$ $=\frac{1}{T}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+\frac{1}{N}\cdot\frac{1}{2}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\mathbb{E}\big{[}\widetilde{Y}_{it}^{2}\big{]}+o\big{(}1/\sqrt{NT}\big{)}$ (2.10)
$=\frac{1}{2}\frac{1}{T}+\frac{1}{2}\frac{1}{N}+o\big{(}1/\sqrt{NT}\big{)}.$
This leads to the conclusion that $\sqrt{NT}\mathbb{E}\big{[}\widehat{l}(\theta)-l(\theta)\big{]}\nrightarrow 0,$ revealing the IPP.
2.3 Bias Correction
The expansion in (2.10) from Example 2 motivates the correction of the likelihood $\widehat{l}(\theta)$ by subtracting the first two terms on the right-hand side of (2.10), ensuring that the corrected likelihood converges to $l(\theta)$ at a rate faster than $1/\sqrt{NT}.$ This paper proposes an analytical bias correction procedure by estimating the bias term $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ from (2.7) under $N/T\rightarrow\kappa\in(0,\infty)$. Specifically, let $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta),$ where $\breve{\alpha}:=(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime})^{\prime}$ and $\breve{\gamma}:=(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime})^{\prime}.$ Accordingly, we partition $[\mathbb{E}H(\theta)]^{-1}$ as
$[\mathbb{E}H(\theta)]^{-1}=:\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
where $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)$ are $(N-1)K\times(N-1)K$ and $(T-1)K\times(T-1)K$ diagonal blocks, respectively, and $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)=[H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)]^{\prime}.$ Lemma 1 (see Section 3) shows that $1/2\mathbb{E}\{s^{\prime}(\theta)[{H}(\theta)]^{-1}s(\theta)\}$ can be asymptotically decomposed into two bias components: $1/2\mathbb{E}\{s_{\breve{\alpha}}^{\prime}(\theta)[H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]s_{\breve{\alpha}}(\theta)\}=:{B}_{\breve{\alpha}}(\theta)$ and $1/2\mathbb{E}\{s_{\breve{\gamma}}^{\prime}(\theta)[H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)]s_{\breve{\gamma}}(\theta)\}=:{B}_{\breve{\gamma}}(\theta),$ resulting from the estimation of the individual and time effects respectively. By estimating the two bias components, we construct the corrected likelihood as follows:
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta).$ (2.11)
The explicit forms of $\widehat{B}_{\breve{\alpha}}(\theta)$ and $\widehat{B}_{\breve{\gamma}}(\theta)$ are defined in Section 3. Subsequently, Theorem 2 demonstrates that
$L(\theta)-l(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
implying that the corrected likelihood no longer suffers from the IPP. Define the MLE of $\theta_{0}$ as the maximizer of $L(\theta):$
$\widehat{\theta}_{L}=\arg\max_{\theta\in\Theta}L(\theta),$ (2.12)
where $\Theta$ is a compact subset of $\mathbb{R}^{K}.$ Under mild regularity conditions, this MLE is asymptotically unbiased, and its asymptotic properties align with the standard results for classical MLE in the independent and identically distributed (i.i.d.) case. Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis.
Since the corrected likelihood depends implicitly on $\widehat{\psi}(\theta),$ which maximizes the reparameterized likelihood function $l^{\ast}(\theta,\psi),$ obtaining $\widehat{\theta}_{L}$ from $L(\theta)$ involves a two-layer optimization. In the inner layer, $\widehat{\psi}(\theta)$ is found by maximizing $l^{\ast}(\theta,\psi)$ using the Newton-Raphson algorithm with analytical derivatives. In the outer layer, we maximize $L(\theta)$ with respect to $\theta,$ substituting $\widehat{\psi}(\theta)$ and using numerical derivatives.
3 Asymptotic Theory
In this section, we develop the asymptotic theory for our bias correction procedure, as well as for the estimation and testing based on the corrected likelihood function. Let $\phi(\theta):=D^{\prime}\psi(\theta)$ and $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta).$ The following assumptions are introduced.
Assumption 1
  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N$ and $T\to\infty.$
  • (ii) Let $\Theta$ be a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior. For each $\theta\in\Theta,$ there exists a compact subset $\Lambda_{\theta}$ that contains both $\widehat{\phi}(\theta)$ and $\phi(\theta)$ in its interior, such that $l(\theta,\phi)$ is three-times continuously differentiable with respect to $\phi\in\Lambda_{\theta}$ almost surely. Let $w_{it}=(Y_{it},X_{it}^{\prime})^{\prime},$ and assume there exists a function $g(w_{it})>0$ independent of $\theta$ and $\phi,$ such that
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Lambda_{\theta}}\left|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ understood as no differentiation), and $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\left[g^{\eta}(w_{it})\right]<\infty,$ for some $\eta>2.$
  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{ik}=\sum_{t=1}^{T}\gamma_{tk}=0$.
  • (iv) Given any fixed $\theta\in\Theta,$ for all $\phi\in\Lambda_{\theta}$ if $\phi\neq\phi(\theta),$ $\mathbb{P}\big{(}l(\theta,\phi)\neq l(\theta,\phi(\theta))\big{)}>0.$
  • (v) Conditional on $\phi(\theta),$$\left\{w_{it}\right\}$ are independent across $i$. Denote $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ to be the $\sigma$-algebras generated, respectively, by $\left\{w_{is}:1\leq s\leq t\right\}$ and $\left\{w_{is}:t\leq s\leq T\right\}$. $\left\{w_{it}\right\}$ is conditional strong mixing with mixing coefficient
    $a_{i}\left(m\right):=\sup_{\theta\in\Theta}\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}\left(A\cap B\mid\phi(\theta)\right)-\mathbb{P}\left(A\mid\phi(\theta)\right)\mathbb{P}\left(B\mid\phi(\theta)\right)\right|,$
    and, $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}\left[a_{i}\left(m\right)\right]^{1-2/\eta}<\infty,$ for some $\eta>2.$
  • (vi) For $\theta\in\Theta,$ $\sqrt{NT}\,\mathbb{E}\left[{H}(\theta)\right]$ is nonsingular when $N$ and $T$ are large.
Assumption 1 requires that both $N$ and $T$ are large and proportional, which is standardis standard in the literature. Assumption 1 imposes smoothness and boundedness conditions on the likelihood. Specifically, it requires that the likelihood and its derivatives for unit $i$ at time $t$ are unformly bounded by a measurable function whose moments of order greater than 2 exist. Assumption 1 is the identification condition for the incidental parameters (fixed effects) as they enter the likelihood additively. Given Assumption 1, Assumption 1 ensures that for each fixed $\theta,$ $\psi(\theta)$ defined in (2.2) is the unique maximizer. Assumption 1 assumes that the data are independent across units $i$ and conditional strong mixing across time $t.$ Under Assumptions 1 and 1, we can apply Davydov’s inequality to control the convergence rate of certain expectations of summations. Assumption 1 requires that the identified incidental parameter Hessian is invertible.
The following lemma states the bias induced by IPP, as shown in (2.7).
Lemma 1
Under Assumption 1, we have
$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\right\}+\mathbb{E}\left\{s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\right\}+o\big{(}1/\sqrt{NT}\big{)},$
where $o(\cdot)$ is uniform in $\theta,$ $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta).$
The following theorem presents the bias terms of $\mathbb{E}[\widehat{l}(\theta)]$ to approximate $\mathbb{E}[l(\theta)].$
Theorem 1
Under Assumption 1, as $N$ and $T\rightarrow\infty,$ we have
$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+B_{\breve{\alpha}}(\theta)+B_{\breve{\gamma}}(\theta)+o\big{(}1/\sqrt{NT}\big{)},$
where $o(\cdot)$ is uniform over $\theta\in\Theta.$ The bias terms $B_{\breve{\alpha}}(\theta)$ and $B_{\breve{\gamma}}(\theta)$ are defined as follows:
$B_{\breve{\alpha}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta){s}_{\breve{\alpha}}(\theta)\big{]},\quad B_{\breve{\gamma}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\gamma}}^{\prime}(\theta){H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta){s}_{\breve{\gamma}}(\theta)\big{]}.$
Since $B_{\breve{\alpha}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ and $B_{\breve{\gamma}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ as $N$ and $T\to\infty,$ these bias terms introduce a non-zero limiting center to the MLE based on $\widehat{l}(\theta).$ Therefore, it is necessary to eliminate the effect of this bias. In the following, we estimate these bias terms and add them back to $\widehat{l}(\theta).$ Let $\psi(\theta)=:(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N-1}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T-1}(\theta))^{\prime},$ $-\sum_{i=1}^{N-1}\alpha_{i}(\theta)=:\alpha_{N}(\theta),$ and $-\sum_{t=1}^{T-1}\gamma_{t}(\theta)=:\gamma_{T}(\theta).$ Then, $\phi(\theta)=(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T}(\theta))^{\prime}.$ Noting that $\mathbb{E}[s_{\breve{\alpha}}(\theta)]=0$ and $\mathbb{E}[s_{\breve{\gamma}}(\theta)]=0,$ we have
$B_{\breve{\alpha}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]^{\prime}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]\big{\}}$
$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}},$ (3.1)
where
$\Sigma^{\breve{\alpha}}(\theta)$ $:=\Sigma_{1}^{\breve{\alpha}}(\theta)+\Sigma_{2}^{\breve{\alpha}}(\theta),$
$\Sigma_{1}^{\breve{\alpha}}(\theta)$ $:=\text{{diag}}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]}\big{\}},$
$\Sigma_{2}^{\breve{\alpha}}(\theta)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\big{[}\widetilde{s}_{N}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},$
$\widetilde{s}_{i}^{\alpha}\left(\theta\right)$ :=$:=\big{\{}\partial_{\alpha_{i}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\alpha_{i}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi\left(\theta\right)},\quad\text{$i=1,\ldots,N.$}$
Similarly, we have
$B_{\breve{\gamma}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]\big{\}}$
$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\gamma}}(\theta)\big{\}},$ (3.2)
where
$\Sigma^{\breve{\gamma}}(\theta)$ $:=\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)][s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}\big{\}}=\mathbb{E}\big{[}\widetilde{S}_{\breve{\gamma}}(\theta)(\widetilde{S}_{\breve{\gamma}}(\theta))^{\prime}\big{]}$
$\widetilde{S}_{\breve{\gamma}}(\theta)$ $:=\big{(}(\widetilde{s}_{1}^{\gamma}(\theta))^{\prime},\ldots,(\widetilde{s}_{T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widetilde{s}_{T}^{\gamma}(\theta)$
$\widetilde{s}_{t}^{\gamma}(\theta)$ :=$:=\big{\{}\partial_{\gamma_{t}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\gamma_{t}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi(\theta)}.$
Next, we estimate the two bias terms. We first estimate the Hessian matrix $H(\theta)=H(\theta,\psi(\theta))$ by substituting $\widehat{\psi}(\theta).$ Define $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta)).$ We have $\widehat{H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\widehat{\phi}(\theta)}\big{]}D^{\prime}$ with $\widehat{\phi}(\theta)=D^{\prime}\widehat{\psi}(\theta).$ Following (3) and (3), we estimate ${B}_{\breve{\alpha}}(\theta)$ and ${B}_{\breve{\gamma}}(\theta)$ by
$\widehat{B}_{\breve{\alpha}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)\right)\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)\right),$
where $\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ are the first $(N-1)K\times(N-1)K$ and last $(T-1)K\times(T-1)K$ blocks of
$[\widehat{H}(\theta)]^{-1}=:\begin{pmatrix}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ \widehat{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
and $\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ and $\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)$ are the estimators of $\Sigma^{\breve{\alpha}}(\theta)$ and $\Sigma^{\breve{\gamma}}(\theta),$ defined as follows. For some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
$\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ $:=\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)+\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta),$
$\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\operatorname*{diag}\big{\{}\widehat{s}_{1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime},\ldots,\widehat{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{\}},$
$\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)$ $:=\left(\iota_{N-1}\iota_{N-1}^{\prime}\right)\otimes\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\widehat{s}_{N,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N,s}^{\alpha}(\theta)\big{)}^{\prime},$
$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=\bigg{[}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)}.$
Here, $\tau$ is a truncation parameter commonly used in the literature (see, e.g., Hahn and Kuersteiner (2011)). When the data exhibit serial dependence, we set $\tau=1$ or $\tau=2$ as in Fernández-Val and Weidner (2016). Our simulations show that the results for $\tau=1$ and $\tau=2$ are similar. For data without serial dependence, we set $\tau=0.$ Additionally, we define
$\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathds{1}\left\{\left|i-j\right|\leq\upsilon\right\}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{j}(\theta)\big{)}^{\prime}$
$\widehat{S}^{\breve{\gamma}}_{i}(\theta)$ $:=\big{(}(\widehat{s}_{i,1}^{\gamma}(\theta))^{\prime},\ldots,(\widehat{s}_{i,T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widehat{s}_{i,T}^{\gamma}(\theta)$
$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=\bigg{[}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)-\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)},\quad i=1,\ldots,N,\ t=1,\ldots,T.$
Specifically, when we set $\upsilon=0$ given conditional cross-sectional independence, $\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ simplifies to
$\widehat{\Sigma}^{\breve{\gamma}}_{0}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{)}^{\prime}.$
Thus, using the estimated bias, we obtain the corrected likelihood $L(\theta):$
$L(\theta)=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta)$
as defined in (2.11).
The following theorem demonstrates that the difference between the corrected likelihood $L(\theta)$ and and the infeasible likelihood $l(\theta)$ is of order $o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$ implying that $L(\theta)$ no longer suffers from IPP.
Theorem 2
Under Assumption 1, the corrected likelihood satisfies
$L(\theta)=l(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
as $N$ and $T\to\infty.$
Compared to $\widehat{l}(\theta),$ the difference between the corrected likelihood and the infeasible likelihood is negligible. The corrected likelihood is derived from the expansion (2.7). Specifically, (2.7) is based on two Taylor expansions: the expansion of the score function for incidental parameters in (2.3) and the expansion of the original likelihood in (2.6). It is seen that the bias terms originate from the quadratic form on the right-hand side of (2.7), expressed in terms of the score and Hessian of the incidental parameters. Under Assumption 1, the quadratic form can be further decomposed into two additive sub-quadratic forms (Lemma 1), capturing the bias contributions from the individual and time dimensions, respectively. Consequently, our bias-corrected likelihood represents the original likelihood augmented by the estimation of these two additive bias terms.
Last, we demonstrate that the MLE obtained from the corrected likelihood, $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta),$ follows an limiting normal distribution with zero mean. Additional regularity conditions are required.
Assumption 2
  • (i) The profile likelihood function $l(\theta)$ is twice differentiable with respect to $\theta\in\Theta,$ where $\Theta$ is a compact subset of $\mathbb{R}^{K}$ with $\theta_{0}$ in its interior.
  • (ii) For all $\theta\in\Theta$ and $\theta\neq\theta_{0},$ $\mathbb{P}\big{(}l(\theta)\neq l(\theta_{0})\big{)}>0.$
  • (iii) For large $N$ and $T,$
    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}|l(\theta)|\big{]}<\infty.$
  • (iv) For large $N$ and $T,$$\{\|\triangledown_{\theta}l(\theta)\|,\theta\in\Theta\}$ is uniformly integrable, and
    $\mathbb{E}\big{[}\sup_{\theta\in\Theta}\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|\big{]}<\infty$
    where $\triangledown_{\theta}$ and $\triangledown_{\theta\theta^{\prime}}$ denote the first and second total derivatives with respect to $\theta.$ Here, $\|\triangledown_{\theta\theta^{\prime}}l(\theta)\|$ defines the Euclidean norm of $\triangledown_{\theta\theta^{\prime}}l(\theta).$
  • (v) The Hessian matrix $\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ is nonsingular for large $N$ and $T$.
Assumption 2 requires the differentiability likelihood on a compact parameter space of $\theta.$ Assumption 2 serves as an identification condition for $\theta$ ensuring that $\mathbb{E}[l(\theta)]$ is uniquely maximized at $\theta_{0}.$ Assumption 2 shows a dominance condition implying that $1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi(\theta))$ converges uniformly in probability to $\mathbb{E}[l(\theta)]$ over $\theta.$ Consequently, Assumptions 2-2 guarantee that $\widehat{\theta}_{L}$ is an asymptotically consistent estimator of $\theta_{0}.$ Assumption 2 establishes a dominance condition for the Hessian, so that for any consistent estimator $\widetilde{\theta}$ of $\theta_{0},$
$1/(NT)\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\widetilde{\theta},\phi(\widetilde{\theta}))\stackrel{{\scriptstyle\mathbb{P}}}{{\longrightarrow}}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})].$
Assumption 2 ensures that the Hessian for $\theta$ is invertible.
Corollary 1
Under Assumptions 1 and 2, we have
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\stackrel{{\scriptstyle\mathbb{D}}}{{\longrightarrow}}\mathcal{N}(0,\Sigma)$
as $N$ and $T\to\infty,$ where $\Sigma=-\lim_{N,T\to\infty}\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1},$ and $\mathcal{N}(0,\Sigma)$ denotes the normal distribution with mean 0 and variance $\Sigma.$
We consider a generic nonlinear null hypothesis $H_{0}:R(\theta_{0})=0,$ where $R(\theta)$ is an $r\times 1$ vector-valued function independent of the data for some $r\leq K.$ Assume that $R(\theta)$ is continuously differentiable, with its Jacobian denoted as $J(\theta):=\triangledown_{\theta^{\prime}}R(\theta),$ and that $\operatorname*{rank}[J(\theta)]=r.$ Let $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0.$ Define the test statistics:
$\widehat{\xi}_{LR}:=-2NT\big{\{}L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\big{\}},$
$\widehat{\xi}_{LM}:=-NT\big{\{}\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\big{\}},$
$\widehat{\xi}_{Wald}:=-NT\big{\{}R(\widehat{\theta}_{L})^{\prime}\big{(}J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\big{)}^{-1}R(\widehat{\theta}_{L})\big{\}}.$
The following corollary states that the hypothesis testing based on the corrected likelihood aligns with standard results in the context of MLE.
Corollary 2
Suppose Assumptions 1 and 2. Under the null hypothesis $H_{0}:R(\theta_{0})=0,$ the statistics $\widehat{\xi}_{LR},$ $\widehat{\xi}_{LM},$ and $\widehat{\xi}_{Wald}$ are asymptotically equivalent, following a common limiting $\chi^{2}(r)$ distribution with degrees of freedom $r$.
4 Simulations
In the simulation study, we consider data generating process (DGP) from a dynamic logit model as follows:
$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ (4.1)
where the disturbance term $\varepsilon_{it}$ follows a standard logistic distribution, independent across both $i$ and $t,$ and also independent of the regressor $X_{it}.$ According to the identification restriction, we generate true values for the incidental parameters $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ from $\mathcal{N}(0,0.01)$, after demeaning. $\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ where we only demean $\gamma_{2,t}^{0}$. Given these parameters, $X_{it}$ is independently generated from $\mathcal{N}\big{(}\sum_{k=1}^{2}(\alpha_{k,i}^{0}+\gamma_{k,t}^{0}),0.25\big{)}$ across both $i$ and $t.$ The true structural parameter is set as $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}.$ The initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ and $\varepsilon_{i0}$ following a standard logistic distribution. Model (4.1) also encompasses the static logit model as a specific case where $\rho\equiv 0.$ We simulate this static case with the same DGP as the dynamic model, except for $\rho_{0}=0$ ($\theta_{0}=\beta_{0}$). We demonstrate the finite sample performance of the MLE of $\theta_{0}$ by maximizing our bias-corrected likelihood $L(\theta),$ and compare it with the uncorrected (original) MLE obtained from $\widehat{l}(\theta)$ without bias correction. In addition, we compare hypothesis testing in terms of size and power based on both the corrected and uncorrected likelihoods. We consider the sample size $(N,T)\in\{(30,30),(60,60),(90,90)\},$ with 1000 replications for each DGP.
Table 1 reports the estimation bias and RMSE of our bias-corrected and the uncorrected MLEs for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models. In the static model, $\widehat{\beta}$ represents the uncorrected MLE from $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is our bias-corrected MLE. For the dynamic model, $(\widehat{\beta},\widehat{\rho})$ denotes the uncorrected MLEs, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected estimates, with the truncation parameter $\tau$ set to $\tau=1$ or $2$. In the static model, ${Y_{it}}$ are generated independently across time $t$ where we set $\tau=0$ for the estimation, while the data in the dynamic model exhibit serial dependence.
Table 1: Comparison between Bias-corrected and Uncorrected MLEs
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.158$ $0.253$ $0.075$ $0.109$ $0.051$ $0.072$
$\widehat{\beta}_{L}$ $0.050$ $0.175$ $0.018$ $0.073$ $0.010$ $0.049$
Dynamic Model
$\widehat{\beta}$ $0.165$ $0.261$ $0.074$ $0.111$ $0.046$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.181$ $0.017$ $0.077$ $0.006$ $0.051$
$\widehat{\beta}_{L}^{(2)}$ $0.061$ $0.183$ $0.018$ $0.078$ $0.007$ $0.051$
$\widehat{\rho}$ $-0.107$ $0.201$ $-0.048$ $0.093$ $-0.033$ $0.058$
$\widehat{\rho}_{L}^{(1)}$ $-0.044$ $0.159$ $-0.014$ $0.076$ $-0.010$ $0.047$
$\widehat{\rho}_{L}^{(2)}$ $-0.038$ $0.161$ $-0.009$ $0.077$ $-0.006$ $0.046$
[…]
    Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model. In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively. In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$. The number of replications is $1000$.
The results in Table 1 reveal that the uncorrected MLEs are severely biased in both models, while our bias-corrected estimators substantially reduce the bias. For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$. As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias for $\widehat{\beta}$ remaining five times larger than that for $\widehat{\beta}_{L}$. Furthermore, the RMSEs form our corrected MLEs are consistently lower than those of the uncorrected MLEs, especially with smaller sample sizes, underscoring the advantage of our bias correction method. Additionally, in the dynamic model, the choice of $\tau$ has little impact on the performance of our bias-corrected estimators.
We also conduct LR tests using both bias-corrected and uncorrected likelihoods. Consider the null hypothesis $H_{0}:\theta_{0}=\theta_{0}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}.$ We report the empirical size ($\delta=0$) and power ($\delta\neq 0$) at a $5\%$ significance level. Table 2 presents the empirical size based on corrected and uncorrected likelihoods, where the corrected likelihoods use truncation $\tau=0$ for the static model and $\tau=1$ for the dynamic model. The results show that the LR tests using uncorrected likelihoods suffer from severe size distortion, while the empirical size from our bias-corrected likelihoods maintains the significance level as the sample size increases. Additionally, our bias-corrected tests significantly reduce Type-I error even for small sample size.
Table 2: Empirical Size of LR Tests Based on Bias-corrected and Uncorrected Likelihoods
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $16\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table 3 presents the empirical power of LR tests based on our bias-corrected likelihoods. The empirical power increases as both the value of $\delta$ and the sample size grow. For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly 100% for both the static and dynamic models. We also perform LM and Wald tests, reporting empirical size and power under the same DGPs in the appendix. The results align with those of the LR tests. Additionally, we perform simulations using probit models with results provided in the appendix.
Table 3: Empirical Power of LR Tests Based on Bias-corrected Likelihoods
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $45\%$ $32\%$ $23\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $63\%$ $37\%$ $25\%$ $49\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $33\%$ $21\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $95\%$ $74\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
    Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Finally, we compare our bias correction method to the Jackknife approach by Fernández-Val and Weidner (2016), who consider a probit model in their simulation. We consider a similar setup, using the same settings as in (4.1), except that the disturbance term $\varepsilon_{it}$ follows a standard normal distribution. Table 4 displays the estimation bias and RMSE of the uncorrected MLEs, our bias-corrected MLEs, and the Jackknife bias-corrected estimators, both for the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) probit models. In the static model, $\widehat{\beta}_{L}$ is our bias-corrected MLE with $\tau=0$, and $\widehat{\beta}_{J}$ represents the Jackknife-corrected MLE. In the dynamic model, $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ is our bias-corrected MLE with $\tau=1$, and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ is the Jackknife-corrected MLE. Table 4 shows that our bias-corrected MLEs outperform both the uncorrected and Jackknife bias-corrected estimators, with consistently smaller RMSEs across all DGPs. The bias correction by our method is significantly greater than that of the Jackknife approach, especially when the sample size is small. For example, in the static model with $(N,T)=(30,30)$, the bias of our corrected MLE is only $0.048$, while the bias of the Jackknife estimator is $-1.228,$ over twenty-five times larger.
Table 4: Comparison with the Jackknife Bias Correction Method
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.144$ $0.197$ $0.061$ $0.081$ $0.041$ $0.052$
$\widehat{\beta}_{L}$ $0.048$ $0.123$ $0.011$ $0.051$ $0.007$ $0.031$
$\widehat{\beta}_{J}$ $-1.228$ $5.930$ $-0.022$ $0.058$ $-0.007$ $0.032$
Dynamic Model
$\widehat{\beta}$ $0.172$ $0.262$ $0.069$ $0.088$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.072$ $0.207$ $0.017$ $0.053$ $0.007$ $0.032$
$\widehat{\beta}_{J}$ $-1.339$ $3.873$ $-0.029$ $0.101$ $-0.009$ $0.034$
$\widehat{\rho}$ $-0.050$ $0.122$ $-0.028$ $0.056$ $-0.017$ $0.037$
$\widehat{\rho}_{L}$ $-0.025$ $0.104$ $-0.014$ $0.049$ $-0.007$ $0.032$
$\widehat{\rho}_{J}$ $0.003$ $0.115$ $-0.001$ $0.050$ $0.000$ $0.033$
[…]
    Notes: $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model. Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected MLE; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators. The number of replications is $1000$.
5 Empirical Analysis
In this section we apply our likelihood-based bias correction method to examining the determinants of the labor force participation (LFP) decision of single mothers. In particular, we look at the impact of the number of children on the decision of the mother to engage in paid employment.
An extensive literature in labor economics has studied the labor supply decisions of married women (Killingsworth and Heckman, 1986; Angrist and Evans, 1998; Blau and Kahn, 2007). These studies have uncovered the impacts of a variety of economic variables, including female market wage and husband income (Mincer, 1962), education (Heath and Jayachandran, 2018), childcare costs (Connelly, 1992), the cost of home technology (Greenwood et al., 2016), and culture norms (Fernández, 2013). Among these variables, the number of children consistently emerges as one of the most important determinants of female labor supply (Nakamura and Nakamura, 1992). This is unsurprising since women continue to bear a disproportionate share of child-rearing responsibilities (Aguero and Marks, 2008).
In contrast to the substantial body of research on the labor supply behavior of married women, the labor supply decisions of single mothers have received relatively limited attention, with only a few exceptions (Kimmel, 1998; Blundell et al., 2016). Single mothers, however, face distinctive challenges when it comes to balancing work and child-rearing responsibilities due to the absence of a second earner in the household. Moreover, their employment decisions may have a more pronounced impact on their children’s well-being compared to the decisions of married women and should thus be of great importance to economists and policymakers.
To study the labor supply decisions of single mothers, we compile a data set from waves $20-30$ of the Panel Study of Income Dynamics (PSID), which span the period of 1987 to 1997. Our sample includes only single mothers, defined as unmarried female household heads with at least one child. The dependent variable is labor force participation status, defined as whether the individual worked any hours during the interview year. Our main explanatory variable is the number of children under 18 in the household. Following Dhaene and Jochmans (2015), we restrict our sample to individuals aged $18-60$ whose LFP status changed at least once during the period. The final sample comprises $N=90$ single mothers whose labor supply status is continuously observed over $T=10$ years from 1987 to 1996. Further details on data construction are provided in Appendix C. On average, 27 percent of these individuals switched in and out of the labor force in each year (Figure C.1--C.2). Table C.1 presents the summary statistics.
Let $Y_{it}\in\left\{0,1\right\}$ denote the labor force participation status of single mother $i$ in year $t$. Let $X_{it}$ denote her number of children. To assess the impact of $X_{it}$ on $Y_{it}$, we estimate the following two logit regressions:
$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\beta X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\}$ (5.1)
$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\left(\beta+\alpha_{i}+\gamma_{t}\right)X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\},$ (5.2)
where $\varepsilon_{it}$ follows a standard logistic distribution, and $\sum_{i}\mu_{i}=\sum_{i}\alpha_{i}=\sum_{i}\gamma_{t}=0$. Model (5.1) is a dynamic homogeneous slope model where $\beta$ represents the average effect of an additional child on the log odds of a single mother’s LFP. Model (5.2) is a dynamic heterogeneous slope model, where $\beta_{it}=\beta+\alpha_{i}+\gamma_{t}$ captures the heterogeneous impact $X_{it}$ may have on the log odds of $Y_{it}$. Model (5.2) nests model (5.1) as a special case when $\beta_{it}$ does not depend on $i$ and $t$.
As discussed in section 1 , when the true impact of $X_{it}$ varies among the population, model (5.1) is appropriate only if $\beta_{it}$ and $X_{it}$ are uncorrelated. In practice, they may correlate due to self-selection. For instance, single mothers who choose to have more children might be those with sufficient financial resources or access to childcare services, such that their labor force participation is less affected by additional children. Alternatively, a correlation between $\beta_{it}$ and $X_{it}$ could arise from common time trends. As we track single mother households from 1987 to 1996, more children were born while factors such as rising childcare and schooling costs, improvements in home technology, and evolving welfare policies changed the impact of each additional child on their mother’s LFP. In both scenarios, model (5.2) would be the appropriate model to use as it controls for unobserved heterogeneity in $\beta_{it}$ by incorporating individual and time fixed effects into the slope and allowing them to correlate with $X_{it}$.
Given our small sample size, direct estimation of both models could suffer from severe incidental parameter problems, which we address using our bias correction procedure. Table 5 presents the estimation results. Panel A reports results for model (5.1) under the homogeneous slope assumption. For each parameter, we provide the MLE and bias-corrected estimates, along with their standard errors. Examining the bias-corrected estimates, we observe that an increase in the number of children appears to reduce a single mother’s likelihood of labor force participation, while past LFP status strongly predicts current LFP. However, this relationship shifts dramatically once we account for slope heterogeneity. Panel B presents the results for model (5.2), allowing for heterogeneous slopes. Here, the bias-corrected estimates reveal a positive average impact of additional children on a mother’s labor supply: each additional child on average increases the log odds of labor force participation by 0.126, which corresponds to an average increase of 5.5 percent in the likelihood of employment. This indicates that, on average, single mothers with more children are indeed more likely to work, likely due to the financial demands of supporting multiple children independently. Additionally, unlike the homogeneous slope model, the serial correlation in labor force participation remains statistically significant but diminishes in magnitude, suggesting that what appears to be strong state dependence in LFP under the homogeneous slope model may in part be driven by unobserved heterogeneity in how children affect labor supply decisions across individuals. Together, these findings underscore the importance of controlling for unobserved slope heterogeneity, as failing to do so can lead to substantially different conclusions.
In addition to highlighting the difference between homogeneous and heterogeneous slope models, Table 5 demonstrates the importance of bias-correction in estimating dynamic nonlinear panel data models. Examining the results in Panel B, we observe that the MLE overestimates the impact of the number of children while underestimating the degree of serial correlation in LFP. Bias correction results in an 18 percent lower estimate of the former and a 57 percent higher estimate of the latter. Moreover, the corrected estimates yield smaller standard errors, improving statistical significance across all parameters.
Finally, in Figure 2, we plot the distribution of the estimated $\beta_{it}$ from model (5.2), with summary statistics in Table C.2. Each $\beta_{it}$ estimate is constructed as the sum of the bias-corrected common effect $\beta$ and the estimated individual and time effects $\alpha_{i}$ and $\gamma_{t}$. These estimates reveal substantial heterogeneity in how the number of children influences single mothers’ labor force participation. While the mean effect is positive, over half of the individual-time observations show a decrease in LFP with each additional child, suggesting that the average effect alone may inadequately capture the diverse impacts across individuals and over time. This reinforces the importance of accounting for both individual and temporal variations in slope coefficients, as the impact of children on labor force participation can vary significantly based on unobserved factors unique to each mother and time period.
In conclusion, our analysis demonstrates the efficacy of our likelihood-based bias correction method for estimating dynamic nonlinear models of labor supply with multidimensional heterogeneities. We show the importance of controlling for unobserved slope heterogeneity when it may correlate with regressors and the value of panel data models that accommodate two-way fixed effects in both the intercept and the slope. Our findings reveal that the number of children has, on average, a positive influence on a single mother’s labor force participation, but such effect is highly heterogeneous across individuals and time periods. This heterogeneity underscores the complex nature of labor supply decisions and the unique challenges each single mother faces in balancing work and child-rearing responsibilities.
Table 5: Single Mother Labor Force Participation: Estimation Results
         Number of Children          Lagged Participation
Uncorrected Corrected Uncorrected Corrected
                Panel A. Homogeneous Slope
Model parameter -0.954 -0.880 0.584 0.535
(0.240) (0.218) (0.113) (0.103)
                Panel B. Heterogeneous Slope
Model parameter 0.153 0.126 0.051 0.080
(0.021) (0.000) (0.020) (0.000)
  • Notes: Standard errors in parentheses. Panel A reports the estimation results of the homogeneous slope model. Panel B reports the estimation results of the heterogeneous slope model. The “uncorrected” and the “corrected” columns show respectively the MLE and the bias-corrected parameter estimates. Data source: PSID 1987 - 1997.
Figure 2: Heterogeneous Impact of Children on Single Mother Labor Force Participation
Acknowledgement
We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu. Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis. Leng’s research is partially supported by the National Natural Science Foundation of China Grant 72001183. Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
Appendix
[…]
AppendixAppendix
A Proofs
A.1 Lemmas
Proof of Lemma 1. Note that ${H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\phi(\theta)}\big{]}D^{\prime},$ where $D$ is a $(N+T-2)K\times(N+T)K$ matix. Since $K$ is fixed, we assume without loss of generality that $K=1.$ We express $H(\theta)$ as:
$H(\theta)=\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&H_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)\end{pmatrix}.$
Consider any element $h_{i,j}^{\breve{\alpha}}(\theta)$ of $N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta),$ where $i,j=1,\ldots,N-1.$ Under Assumptions 1 and 1, by applying Davydov’s inequality, it follows that
$\sup_{i,j}\sup_{\theta\in\Theta}\text{Var}\big{(}h_{i,j}^{\breve{\alpha}}(\theta)\big{)}=o(1)$
as $N$ and $T\rightarrow\infty.$ Consequently, by the Markov inequality,
$\sup_{i,j}\sup_{\theta\in\Theta}\big{\{}h_{i,j}^{\breve{\alpha}}(\theta)-\mathbb{E}[h_{i,j}^{\breve{\alpha}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Let $h_{t,s}^{\breve{\gamma}}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)$ for $t,s=1,\ldots,T-1.$ By conditional cross-sectional independence, we also have
$\sup_{t,s}\sup_{\theta\in\Theta}\big{\{}h_{t,s}^{\breve{\gamma}}(\theta)-\mathbb{E}[h_{t,s}^{\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Moreover, for each $h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta),$ the $(i,t)$-th element of $NT\cdot H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta),$ with $i=1,\ldots,N-1$ and $t=1,\ldots,T-1,$ by Assumption 1 we have that
$\sup_{i,t}\sup_{\theta\in\Theta}\big{\{}h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)-\mathbb{E}[h_{i,t}^{\breve{\alpha}\breve{\gamma}}(\theta)]\big{\}}=o_{\mathbb{P}}(1).$
Therefore,
$N\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),\quad T\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
and
$NT\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)-\mathbb{E}[H_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
uniformly for $\theta\in\Theta,$ where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. Since $\mathbb{E}[s(\theta)]=0,$ $s(\theta)=s(\theta)-\mathbb{E}[s(\theta)]=:(\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta),\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta))^{\prime}.$ We now decompose the quadratic form on the right hand side of (2.7):
$s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
$=$ $[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]$
$=$ $[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-\mathbb{E}{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]^{\prime}(1+o_{\mathbb{P}}(1))$
$=$ $-\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)+2\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}\left(\theta\right)\big{]}(1+o_{\mathbb{P}}(1)).$ (A.1)
By matrix inversion in block form, it follows that
$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(1),$
and
$\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\right\|_{\max}=O(T),\quad\sup_{\theta\in\Theta}\left\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N).$
In the following, we show
$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O\big{(}1/(NT)\big{)},$ (A.2)
and
$\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}=O(1/T),\quad\sup_{\theta\in\Theta}\,\mathbb{E}\big{[}\widetilde{s}_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}=O(1/N).$ (A.3)
Therefore, $\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right)H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)\big{]}$ is negligible in the expansion for $\mathbb{E}[\widehat{l}(\theta)]$ compared to the other two terms from the quadratic form.
Let
$\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
and
$\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta)):=\{\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)-\mathbb{E}[\partial_{\gamma_{t}}l_{i,t}(\theta,\phi)]\}|_{\phi=\phi(\theta)}$
for $i=1,\ldots,N$ and $t=1,\ldots,T,$ respectively. Define $h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)$ as the $(i,t)$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ where $i=1,\ldots,N-1$ and $t=1,\ldots,T-1.$ Under the assumption of conditional cross-sectional independence, we have
$\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}\left(\theta\right))H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\breve{\gamma}}(\theta))\big{\}}$
$=$ $\frac{1}{\sqrt{NT}}\mathbb{E}\bigg{\{}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\cdot\bigg{[}\sum_{s=1}^{T}\big{(}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{)}\bigg{]}$
$\hskip 120.00018pt\cdot\bigg{[}\sum_{j=1}^{N}\big{(}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{)}\bigg{]}\bigg{\}}$
$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right]$
$\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right]$
$\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\bigg{)}\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]$
$\quad+\frac{1}{\sqrt{NT}}\bigg{(}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{N,T}(\theta,\phi(\theta))\bigg{)}$
=:absent:=:= : $A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
By Assumptions 1 and 1, using Davydov’s inequality, we obtain for some $\eta>2$ that
$|A_{NT,1}(\theta)|$
$=$ $\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\bigg{(}\sum_{s=1}^{T-1}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\bigg{)}\bigg{(}\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right.\right.$
$\hskip 56.9055pt\left.\left.+\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\cdot h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\right]\right|$
$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O\big{(}\sqrt{NT}\big{)},$
$|A_{NT,2}(\theta)|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\bigg{(}\frac{1}{T}\sum_{t=1}^{T-1}|h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)|\bigg{)}\cdot\left|\mathbb{E}\bigg{(}\sum_{s=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\partial_{\gamma_{T}}\widetilde{l}_{i,T}(\theta,\phi(\theta))\bigg{)}\right|$
$\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)},$
$|A_{NT,3}(\theta)|$
$=$ $\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\right)\left(\sum_{t=1}^{T-1}\partial_{\gamma_{t}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\right)\right]\right|$
$\leq$ $\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}$
$\quad\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\big{(}1-m/(T-1)\big{)}[a_{i}(m)]^{1-2/\eta}+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O\big{(}\sqrt{NT}\big{)},$
and
$|A_{NT,4}(\theta)|\leq\sqrt{NT}\sup_{\theta\in\Theta}\left\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\right\|_{\max}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)=O\big{(}\sqrt{NT}\big{)}.$
Thus, we obtain (A.2) from (A.4).
Let $h_{i,j}^{\ast\breve{\alpha}}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta),$ where $i,j=1,\ldots,N-1.$ By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ uniformly in $\theta.$ Define $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}.$ Noting that
$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
$N\cdot H_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix. Thus, it follows from the lemma in Miller (1981) that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Since $s(\theta)=D\partial_{\phi}l(\theta,\phi)|_{\phi=\phi(\theta)},$ we have
$T\mathbb{E}\big{[}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\big{]}$
$=$ $\frac{1}{N}\mathbb{E}\big{\{}(N\sqrt{T}\widetilde{s}_{\breve{\alpha}}^{\prime}(\theta))[N^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\breve{\alpha}}(\theta))\big{\}}$
$=$ $\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}^{2}$
$+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\,\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}\right.$
$\hskip 116.65646pt\left.\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{(}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{)}\bigg{)}\right]$
=:absent:=:= : $I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
We now show that $I_{NT,1}(\theta)=O(1)$ uniformly for $\theta.$ By Assumptions 1 and 1, applying Davydov’s inequality gives that for some $\eta>2,$
$\sup_{\theta\in\Theta}|I_{NT,1}(\theta)|$
$=$ $\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|$
$\hskip 42.67912pt\cdot\bigg{|}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))\partial_{\alpha_{i}}\widetilde{l}_{i,s}(\theta,\phi(\theta))\big{]}+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\big{[}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\partial_{\alpha_{N}}\widetilde{l}_{N,s}(\theta,\phi(\theta))\big{]}\bigg{|}$
$\leq\bigg{(}\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
$=O(1)$
as $N$ and $T\rightarrow\infty,$ where the conditional cross-sectional independence leads to the first equality. Similarly, for $I_{NT,2}(\theta),$ we have
$\sup_{\theta\in\Theta}|I_{NT,2}(\theta)|$
$=$ $\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\cdot\mathbb{E}\left[\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}^{2}\right]$
$\leq$ $\bigg{(}\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\breve{\alpha}}(\theta)|\bigg{)}\cdot\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}(1-m/T)\cdot[a_{i}(m)]^{1-2/\eta}\right)$
$=$ $O(1),$
noting that $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta.$ Thus, from (A.5), we obtain the first result in (A.3). Similarly, we can derive the other result in (A.3). Now, combining (A.1), (A.2), and (A.3) yields that
$\mathbb{E}\big{\{}s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\big{\}}=$ $-\mathbb{E}\big{\{}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\big{\}}-\mathbb{E}\big{\{}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\big{\}}$
$-\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}.$
Note that, by Assumption 1, for $S\in\{0,1,2\},$
$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
as $a\rightarrow\infty,$ which shows that $\{\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)\}_{i,t}$ are uniformly integrable. By the $L_{1}$ Convergence Theorem, we hence have
$\sqrt{NT}\mathbb{E}\left\{\big{[}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)\big{]}\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
as $N,T\rightarrow\infty.$ This completes the proof of Lemma 1.
Lemma A.1
Under Assumption 1, $r^{s}(\theta,a)$ defined by (2.3) satisfies
$\sup_{\theta\in\Theta}\sup_{a}\left\|r^{s}(\theta,a)\right\|_{\max}=O_{\mathbb{P}}\left(1/\left(NT\right)\right),$
as $N$ and $T\rightarrow\infty$, where $a=\left(a_{1},\ldots,a_{\left(N+T-2\right)K}\right)^{\prime}$ for each component $a_{j}\in(0,1)$.
Proof. In the proof, we retain the notation and setup from Lemma 1. From equations (2.3) and (2.5), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta).$ Specifically,
$[-{H}(\theta)]^{-1}s(\theta)$ $=[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)](1+o_{\mathbb{P}}(1))$
$=-\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.6)
where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t.$ Let $u_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)$ for $i=1,\ldots,N-1.$ It follows that
$N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\big{[}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.7)
Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\breve{\alpha}}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in $\theta,$ and it also implies that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta.$ Since $\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{i}}\widetilde{l}_{i,t}(\theta,\phi(\theta))=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{j}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1)$
and
$N\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot\bigg{(}\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\partial_{\alpha_{N}}\widetilde{l}_{N,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1).$
Therefore, $N\sqrt{T}u_{i}^{\breve{\alpha}}(\theta)=O_{\mathbb{P}}(1)$ uniformly for $i.$ Let $u_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Given that $\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\|_{\max}=O(1),$ we have
$\sqrt{N}Tu_{i}^{\breve{\alpha}\breve{\gamma}}(\theta)=\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\big{[}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))-\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\big{]}\bigg{)}.$ (A.8)
Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)=O(1),$ which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}l^{*}(\theta,\psi)|_{\psi=\psi(\theta)}]=0,$ it holds that
$\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{t}}\widetilde{l}_{j,t}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
and
$\bigg{(}\sum_{t=1}^{T-1}h_{i,t}^{\ast\breve{\alpha}\breve{\gamma}}(\theta)\bigg{)}\bigg{(}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\partial_{\gamma_{T}}\widetilde{l}_{j,T}(\theta,\phi(\theta))\bigg{)}=O_{\mathbb{P}}(1),$
uniformly in $\theta.$ From (A.7) and (A.8), we conclude that
$\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)}.$ (A.9)
Similarly, we can show that
$\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(\sqrt{N}T)\big{)},\quad\sup_{\theta\in\Theta}\|H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\big{(}1/(N\sqrt{T})\big{)}.$ (A.10)
Thus, combining (A.6), (A.9), and (A.10), we have
$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)},$
where $o_{\mathbb{P}}(\cdot)$ is uniform in $\theta.$ Noting that
$\widehat{\psi}(\theta)-\psi(\theta)=[-{H}\left(\theta\right)]^{-1}s(\theta)(1+o_{\mathbb{P}}(1)),$
we further have
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\big{(}1/(NT)^{3/4}\big{)}.$ (A.11)
From (A.11), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1.$ We can express $\partial_{\alpha_{i}}H(\theta)$ as
$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\end{pmatrix}.$
In particular, let $\text{diag}\{a_{1},\ldots,a_{n}\}$ denote a diagonal matrix with elements $a_{1},\ldots,a_{n}$ on its diagonal. We then have
$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)=$ $\frac{1}{T}\sum_{t=1}^{T}\text{diag}\big{\{}0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\big{\}}$
$\quad-\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\text{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
$\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)=$ $\big{(}\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes e_{i}$
$\quad+\big{(}\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\big{)}\otimes\iota_{N-1},$
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to 1 and all other elements equal to 0. Using (A.11), for each $j=1,\ldots,N-1,$ we write
$\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $\sup_{a_{j}\in(0,1)}\frac{1}{2}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}G_{\psi_{j}}(\theta,a_{j})\big{(}\widehat{\psi}(\theta)-\psi\left(\theta\right)\big{)}$
$=$ $\frac{1}{2}\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\\ H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\end{pmatrix}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
$=$ $\frac{1}{2}\left\{\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}(\theta)\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right.$
$\qquad+\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}$
$\qquad\left.+2\big{(}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\big{(}H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)\big{)}\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}$
=:absent:=:= : $\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\big{(}1+o_{\mathbb{P}}(1)\big{)}.$ (A.12)
Let $\widetilde{u}_{i}^{\breve{\alpha}}(\theta)$ denote the $i$-th element of $H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $i=1,\ldots,N-1.$ Using (A.7) and (A.8), we can show that
$C_{NT,1}(\theta)$
$=$ $\frac{1}{N}[\widetilde{u}_{i}^{\breve{\alpha}}(\theta)]^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}-\frac{1}{N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
uniformly in $\theta.$ By similar arguments, we also show that
$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)},\quad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\big{(}1/(NT)^{3/2}\big{)}.$
Therefore, when $\psi_{j}=\alpha_{i},$ we have
$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $-\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.13)
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)}.$
Similarly, when $\psi_{j}=\gamma_{t},$ we have
$\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=$ $-\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}+o_{\mathbb{P}}\big{(}1/(NT)\big{)}$ (A.14)
$=$ $O_{\mathbb{P}}\big{(}1/(NT)\big{)},$
where $\widetilde{u}_{t}^{\breve{\gamma}}(\theta)$ denotes the $t$-th element of $H_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)+H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\gamma}}(\theta)$ for $t=1,\ldots,T-1.$ This shows the desired result.
Lemma A.2
Under Assumption 1, $r(\theta)$ defined by (2.7) satisfies
$\sup_{\theta\in\Theta}\mathbb{E}[r(\theta)]=O\left(1/\left(NT\right)\right),$
as $N$ and $T\rightarrow\infty$.
Proof. In the proof, we retain the notation and setup from Lemmas 1 and A.1. From Lemma A.1, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta).$ Following similar arguments as in Lemma 1, we can show that
$\mathbb{E}\big{[}(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\big{]}=O\big{(}1/(NT)\big{)}.$
Moreover, from (A.11), (A.13), and (A.14), we have
$\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
$=$ $\frac{1}{2N}\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$+\frac{1}{2T}\bigg{(}\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}^{2}\bigg{(}\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\bigg{)}$
$\quad+\bigg{(}\sum_{i=1}^{N-1}\widetilde{u}_{i}^{\breve{\alpha}}(\theta)+\sum_{t=1}^{T-1}\widetilde{u}_{t}^{\breve{\gamma}}(\theta)\bigg{)}\cdot o_{\mathbb{P}}\big{(}1/(NT)\big{)}$
$=$ $O_{\mathbb{P}}\big{(}1/(NT)^{5/4}\big{)}.$ (A.15)
By Assumption 1 and applying the $L_{1}$ Convergence Theorem to (A.15), we find that
$\sup_{\theta\in\Theta}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=o\big{(}1/(NT)\big{)}.$
Thus,
$\mathbb{E}\big{[}r(\theta)\big{]}=\frac{1}{2}\mathbb{E}\big{[}(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\big{]}+\frac{1}{3}\mathbb{E}\big{[}\big{(}\widehat{\psi}(\theta)-\psi(\theta)\big{)}^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\big{]}=O\big{(}1/(NT)\big{)}$
uniformly in $\theta.$
A.2 Proof of Theorem 1
From (2.7), we have
$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+\frac{1}{2}\mathbb{E}\big{\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\big{\}}-\mathbb{E}[r(\theta)].$
Thus, the desired result follows directly from Lemmas 1 and A.2.
A.3 Proof of Theorem 2
We retain the notation introduced in Lemmas 1-A.2. First, we show that
$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.16)
Note that
$\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=\frac{1}{2}\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}+\frac{1}{2}\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}.$
Define $\widetilde{s}_{i,t}^{\alpha}(\theta):=\{\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})-\mathbb{E}[\partial_{\alpha_{i}}l_{i,t}(\theta,{\phi})]\}|_{\phi={\phi}(\theta)}.$ Then we have
$N^{2}\big{[}{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\big{]}$
$=$ $\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}$
[…]
$+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
×$\qquad\times\text{diag}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}\big{\}}.$ (A.17)
Under Assumptions 1 and 1, by Davydov’s inequality,
$\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}=\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\big{]}\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
where we assume without loss of generality that $K=1.$ Thus, as $T\rightarrow\infty,$
$\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\big{[}\widetilde{s}_{i,t}^{\alpha}(\theta)\big{(}\widetilde{s}_{i,s}^{\alpha}(\theta)\big{)}^{\prime}\big{]}$
$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}$
$\leq$ $8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\cdot\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}$
$=$ $o(1)$
given that $\tau/T\to 0.$ Combining this with (A.17) implies
$\sup_{\theta\in\Theta}N^{2}\|{\Sigma}_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.18)
From the proof of Lemma A.1, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1).$ Moreover, under the conditional strong mixing condition, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to its counterpart in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore, from (A.18),
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{1}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma^{\breve{\alpha}}_{1}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{1,\tau}(\theta)]\big{\}}$
$=$ $\frac{T}{N}\cdot\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
$=$ $o_{\mathbb{P}}(1/T).$
Similarly,
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\text{trace}\big{\{}T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\,N^{2}[\Sigma_{2}^{\breve{\alpha}}(\theta)-\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)]\big{\}}$
$=$ $\frac{T}{N^{2}}\cdot\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\cdot o_{\mathbb{P}}(1/T)$
$=$ $o_{\mathbb{P}}(1/N^{2}),$
where the last equality follows from the fact that $\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta,$ as shown in Lemma A.1. Therefore,
$\text{trace}\big{\{}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)[\Sigma^{\breve{\alpha}}(\theta)-\widehat{\Sigma}^{\breve{\alpha}}_{\tau}(\theta)]\big{\}}=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.19)
Moreover, from Theorem 1, we also have
$\text{trace}\big{\{}[\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)-H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)]\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot\text{trace}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}}=o_{\mathbb{P}}(1)\cdot B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$ (A.20)
Combining (A.19) and (A.20) shows the first part of (A.16). The second part of (A.16) can be shown similarly.
Noting that
$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)+\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$
as implied by the proof of Theorem 1, we have
$L(\theta)-l(\theta)$ $=\big{[}\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\big{]}+\big{[}\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\big{]}$
$\quad-\bigg{[}\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\bigg{]}-\bigg{[}\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\bigg{]}$
$+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
Moreover, from Lemma 1,
$\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\frac{1}{2}s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
Thus, combining this with (A.16) yields the desired result.
A.4 Proof of Corollary 1
Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}:$
$0=\triangledown_{\theta}L(\widehat{\theta}_{L})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$ (A.21)
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\theta_{0}.$ From Theorem 2, we further have
$0=$ $\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$ (A.22)
Under Assumption 2, $\theta_{0}$ is the unique maximizer of $\mathbb{E}[l(\theta)].$ Since $L(\theta)-\mathbb{E}[l(\theta)]=o_{\mathbb{P}}(1),$ under Assumption 2, we conclude that
$\widehat{\theta}_{L}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}.$
This also implies $\overline{\theta}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}.$ By Assumption 2, it follows that
$\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]=o_{\mathbb{P}}(1).$
Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (A.22) that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})$ $=-\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.23)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.23) shows that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\Sigma^{-1}\Omega\Sigma^{-1}\right)$ (A.24)
where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0,$ and
$\Omega=\lim_{N,T\rightarrow\infty}(NT)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
Under Assumption 2, the information matrix equality ($\Sigma^{-1}=\Omega$) holds, where the asymptotic variance in (A.24) simplifies to $\Sigma$.
A.5 Proof of Lemma A.3
Lemma A.3
Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\to\infty,$
$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where
$\Xi(\theta_{0}):=\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{(}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{)}^{-1}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}.$
Proof. Consider the constrained maximization problem:
$\big{(}\widehat{\theta}_{R},\widehat{\lambda}\big{)}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\big{[}L(\theta)+\lambda^{\prime}R(\theta)\big{]}$
where $\lambda$ is the Lagrangian multiplier. The first-order conditions for this constrained maximization are
$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (A.25)
$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (A.26)
Following a similar proof as in Corollary 1, we have $\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=O_{\mathbb{P}}\left(1\right).$ Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.25) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, we express the mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ as
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}$
$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (A.25) yields that
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.27)
Under $H_{0},$ from (A.26), we also obtain
00 0 $=\sqrt{NT}R(\widehat{\theta}_{R})$
$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1)$
$=J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1).$ (A.28)
Now, combining (A.27) and (A.28), we have
$\begin{pmatrix}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]&\left[J(\theta_{0})\right]^{\prime}\\ J(\theta_{0})&0\end{pmatrix}\begin{pmatrix}\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})\\ \sqrt{NT}\widehat{\lambda}\end{pmatrix}=\begin{pmatrix}-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\\ 0\end{pmatrix}+o_{\mathbb{P}}(1)$
which has the solution
$\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=-\left(\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.29)
The conclusion follows by subtracting (A.23) in Corollary 1 from both sides of (A.29).
A.6 Proof of Corollary 2
Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0}).$ We first show that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (A.30)
A mean value expansion of $L(\widehat{\theta}_{R})$ around $\widehat{\theta}_{L}$ gives
${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+\frac{1}{2}\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
where each component of $\theta^{\ast}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Noting that $\triangledown_{\theta}{L}(\widehat{\theta}_{L})=0,$ we obtain
$\widehat{\xi}_{LR}=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}.$
Since $\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1),$ by Lemma A.3, we further conclude that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1).$
For $\widehat{\xi}_{LM},$ note that
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}.$ Thus,
$\widehat{\xi}_{LM}$ $=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
$=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
$=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
For $\widehat{\xi}_{Wald}$, a mean value expansion of $R(\widehat{\theta}_{L})$ gives
$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
[…]
$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\ddot{\theta})^{\prime}\left(J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right)^{-1}J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
$=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
Next we show
$\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right).$ (A.31)
Since $\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\Sigma),$ by the delta method, we have
$J(\theta_{0})\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\big{(}0,J(\theta_{0})\Sigma J(\theta_{0})^{\prime}\big{)}.$
This implies (A.31). Therefore, the desired result follows from (A.30) and (A.31). The proof is complete.
B Additional Simulation Results
B.1 Logit Model
Table B.1: Empirical Size of LM and Wald Tests for Logit Model
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$
(a) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $32\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
(b) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $47\%$ $31\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $39\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
B.2 Probit Model
Table B.3: Comparison between the Bias-corrected and Uncorrected MLE
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.150$ $0.197$ $0.063$ $0.082$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.055$ $0.124$ $0.013$ $0.049$ $0.007$ $0.032$
Dynamic Model
$\widehat{\beta}$ $0.159$ $0.247$ $0.067$ $0.086$ $0.043$ $0.055$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.134$ $0.015$ $0.052$ $0.008$ $0.032$
$\widehat{\beta}_{L}^{(2)}$ $0.062$ $0.136$ $0.016$ $0.053$ $0.008$ $0.032$
$\widehat{\rho}$ $-0.055$ $0.124$ $-0.022$ $0.054$ $-0.017$ $0.036$
$\widehat{\rho}_{L}^{(1)}$ $-0.030$ $0.105$ $-0.009$ $0.047$ $-0.007$ $0.031$
$\widehat{\rho}_{L}^{(2)}$ $-0.022$ $0.106$ $-0.003$ $0.047$ $-0.003$ $0.031$
[…]
Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model. In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively. In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$. The number of replications is $1000$.
Table B.4: Empirical Size of LR, LM, and Wald Tests Based on Bias-corrected and Uncorrected Likelihoods for Probit Model
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $29\%$ $13\%$ $27\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $23\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $26\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $25\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.5: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $41\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $96\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $63\%$ $40\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $58\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $95\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $82\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
[…]
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
C Additional Empirical Results
C.1 Sample
The data set is constructed from waves $20-30$ of PSID, covering the survey years 1987 to 1997. We selected this time span because 1997 was the last year PSID conducted annual surveys, switching to a biennial schedule in 1999. Each survey wave includes questions about employment status in the prior calendar year, such as, ‘About how many weeks did (he/she) work on that job last year?’ and ‘During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?’ We use responses to these questions to construct individual labor force participation (LFP) status for the prior year. Thus, wave 30 (1997) provides LFP data for 1996, wave 29 (1996) for 1995, and so on. In our final sample, labor force participation status is observed from 1987 to 1996, with lagged participation status observed from 1986 to 1995.
We also restrict our sample to individuals whose LFP status changed at least once during the sample period. Individuals whose labor supply status never changed during the sample period are not informative for our analysis, because the dependent variable could be perfectly predicted by setting the individual fixed effects in the intercept of our models to plus or negative infinity.
The final sample consists of a panel of $N=90$ individuals over $T=10$ years. Table C.1 presents the summary statistics. Figure C.1 shows the average percentage of individuals with given current and past LFP status. Figure C.2 plots the LFP transition rates by year.
Table C.1: Descriptive Statistics
Variable Mean Std. Dev. Min Max
Labor Force Participation 0.482 0.500 0.000 1.000
Lagged Participation 0.458 0.498 0.000 1.000
Number of Children 2.24 1.098 1.000 6.000
  • Notes: Summary statistics are computed based on a sample of 90 individuals observed from 1987 to 1996. All variables have 900 observations. Data source: PSID 1987 - 1997.
Figure C.1: Labor Force Participation Contingency Table. Each cell contains the percentage of individuals with given current and past labor force participation status.
Figure C.2: Labor Force Participation Status Transition Rates By Year
C.2 Extra Empirical Results
Table C.2: Estimated Heterogeneous Effects on Labor Force Participation
10th 25th 75th 90th
Mean Median Std. Dev. Percentile Percentile Percentile Percentile
$\alpha_{i}$ $0$0.975 6.203 -6.909 -2.443 4.292 6.774
$\gamma_{t}$ 0.000 0.146 0.773 -1.151 -0.466 0.638 0.979
$\beta_{it}$ 0.126 -0.504 6.250 -6.835 -2.628 4.050 6.982