### v4

# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time fixed effects in both the intercept and slope. These models are subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at zero, and that test statistics do not follow their standard asymptotic distributions as in the absence of the fixed effects. To address the problem, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting estimator follows an asymptotic normal distribution with mean zero. Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
Keywords:
incidental parameter problem; bias correction; fixed effects; panel data models
JEL Classification:
C23
1 Introduction
Panel data is common in empirical research. Panel data models with two-way fixed effects are widely used to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007). When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
To illustrate, consider the classic problem of estimating the impact of children on a mother’s labor force participation. Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins 1988; Angrist and Evans 1998; Kleven et al. 2019). First, mothers who choose to have more children might be inherently less inclined to work. Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or access to childcare. In a panel dataset tracking households over time, we can control for the first type of level heterogeneity by including fixed effects in the intercept. However, to address the second type of response heterogeneity, we must include fixed effects in the slope. Given that labor force participation is a binary outcome and typically exhibits serial correlation, this example underscores the need for a dynamic nonlinear panel data model with fixed effects in both the intercept and slope.
In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models. DN-HP models are generally subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at 00, and that the test statistics do not follow their standard asymptotic distributions (of $\chi^{2}$, for many test statistics). To address the incidental parameter problem, we propose an analytical bias correction approach. This method restores the standard asymptotic distributions for the point estimator and the test statistics.
Figure 1: Comparison of Uncorrected and Corrected Maximum Likelihood Estimators
{justify}
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$. $\widehat{\beta}$ and $\widehat{\rho}$ are the uncorrected estimators while $\widehat{\beta}_{L}$ and $\widehat{\rho}_{L}$ are the corrected estimators. $1000$ replications. The detailed data-generating process is given in Section 4.
As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing estimators from our bias correction procedure to the uncorrected ones. The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean square errors.
Literature Review. The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity. Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept. To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework. For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator. Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods. Woutersen (2002) and Li et al. (2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood. Most of these methods are analytical, estimating the bias using explicit formulas. Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of $O(1/N)$. Moon and Weidner (2017) propose a parameter-based analytical correction for dynamic linear models with two-way interactive effects. For nonlinear models, Fernández-Val and Weidner (2016) develop parameter-based bias correction techniques with additive fixed effects, while Chen et al. (2021) focus on interactive fixed effects in static models. Alternatively, Jochmans and Otsu (2019) propose a likelihood-based analytical bias correction for static models, allowing fixed effects to enter the model in an arbitrary but known form. Other model-specific solutions include Bai (2009) and Charbonneau (2017), among others. For an extensive review of the incidental parameter problem for models with two-way fixed effects in the intercept, see Fernández-Val and Weidner (2018).
To our knowledge, our paper represents one of the first efforts to extend intercept-only two-way fixed effects models into multi-dimensional two-way fixed effects models, accounting for both level and response heterogeneities. Two prior studies in this line of research are Keane and Neal (2020) and Lu and Su (2023). Both examine linear heterogeneous parameter models with additive effects in both the intercept and slope. Keane and Neal (2020) propose an iterative procedure, which they refer to as the “mean observation OLS,” to eliminate the bias, while Lu and Su (2023) rely on a parameter-based Jackknife approach and provide a uniform inference theory for structural parameters. Our study significantly expands the model framework and introduces a likelihood-based analytical bias correction technique to ensure valid estimation and inference. Neither our DN-HP models nor our bias correction procedure have been discussed in the current literature.
This paper. We employ the maximum likelihood framework and focus on multi-dimensional additive two-way fixed effects, whose number grows with the sample size. We consider an arbitrary DN-HP model whose log-likelihood function is specified up to the unknown parameters. We construct a modified, or corrected, log-likelihood function by adding two bias correction terms to the original one. These two bias terms are analytically derived by combining i) the Taylor expansion of the original log-likelihood function, in terms of the fixed effects, and ii) an asymptotic expansion of the fixed effects themselves. We show that the corrected log-likelihood function does not suffer from the incidental parameter problem, under appropriate regularity conditions.
One key benefit of our procedure is that we achieve bias corrections for the point estimators and the test statistics by modifying a single object, the log-likelihood function. We rigorously show that the estimators obtained by maximizing the corrected log-likelihood function retains a limiting normal distribution with zero mean, consistent with the asymptotic theory of the classical maximum likelihood estimators (MLE) in the absence of incidental parameters. Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife, which is a potential bias correction device [^1] It is worth mentioning that, to our knowledge, there currently has been no study on using the jackknife under the DN-HP models.. We show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors, and restores the test sizes of the LR test to its nominal level. We also find that our procedure outperforms the Jackknife. In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
The remainder of the paper is organized as follows. Section 2 explains our settings, the identification restriction, and the incidental parameter problem of DN-HP models. Section 3 describes our bias correction procedure and provides relevant statistical properties, including those we claim above. Section 4 presents simulation studies to demonstrate the performance of method. Section 5 applies our method to study single mothers’ labor force participation. Finally in Section 6, we leave some closing remarks. Proofs, technical details, and additional results are provided in the appendix.
Notation. We denote $\mathbb{I}_{n}$ to be the $n\times n$ identity matrix and $\iota_{n}$ the $n\times 1$ vector of ones. $\otimes$ denotes the Kronecker product and $\mathds{1}\{\cdot\}$ is the indicator function. Next, for a sequence of square matrices $A_{i}$, $i=1,\ldots,n$, $\operatorname*{diag}\{A_{1},\ldots,A_{n}\}$ represents a block diagonal matrix where each $A_{i}$ is the $i$-th diagonal element. For a vector $v$ and a function $f(v)$, we write $\partial_{v}$ and $\partial_{vv^{\prime}}$ to represent the first and second partial derivatives of $f(v)$ with respect to (w.r.t.) $v$.
2 Bias in Heterogeneous Parameter Models
2.1 Models and Estimations
Let $\{(Y_{it},X_{it}^{\prime})^{\prime}:i=1,\ldots,N;t=1,\ldots,T\}$ be a panel data set of $N$ individuals and $T$ time periods, where, $i$ represents the individual and $t$ represents the time period. Here $Y_{it}$ is a scalar response variable, and $X_{it}$ is a $K\times 1$ vector of regressors, with $K$ known and fixed. For each individual $i$, the response $Y_{it}$ is generated sequentially over $t$, by
$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)
where $\theta_{0}$ is a $K\times 1$ vector of (non-random) unknown common parameters of interests, $\phi^{0}:=(\alpha_{1}^{0\prime},\ldots,\alpha_{N}^{0\prime},\gamma_{1}^{0\prime},\ldots,\gamma_{T}^{0\prime})^{\prime}$ consists of the $K\times 1$ vectors of random variables $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, and $f(y|\cdot)$ is a density known up to $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$. We restrict our attention to the case where $f\left(\cdot\right)$ depends on $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$ through the additive structure $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, the “heterogeneous parameter”. The vectors $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ capture the individual-specific and time-specific heterogeneities, respectively. Generally, $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ may be correlated with the regressor $X_{it}$. However, throughout the paper, we employ the “fixed-effect” framework to condition on the realizations of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ (denoted by the same symbols) in the density, and treat them as unknown nuisance parameters to be estimated together with $\theta_{0}$. Conditioning on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, we assume $(Y_{it},X_{it}^{\prime})^{\prime}$ independent across $i$ but serially dependent over $t$. In particular, we allow $X_{it}$ to contain both strictly exogenous components and predetermined components w.r.t. $Y_{it}$. For instance, $X_{it}$ may contain lagged values of $Y_{it}$. In this paper, we focus on the case where only heterogeneous parameters appear in (2.1). This is without loss of generality. As illustrated in the example below, our approach accommodates settings where both heterogeneous parameters and “homogeneous parameters” are present. [^2] Another example is the classic Neyman-Scott model, where $Y_{it}$ follows a normal distribution with a homogeneous variance but heterogeneous means. See Example 3 below.
Example 1 (heterogeneous parameter model)
Consider a probit model with data generating process (DGP)
$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}=\mathds{1}\{\pi(X_{it},\theta_{0},\alpha_{i}^{0},\gamma_{t}^{0})+\varepsilon_{it}>0\},$ (2.2)
where $\varepsilon_{it}$ is a standard normal random variable independent of $X_{it}$, $\theta_{0}$, $\alpha_{i}^{0}$, and $\gamma_{t}^{0}$. For arguments $\theta$, $\alpha_{i}$, and $\gamma_{t}$ of dimension $K\times 1$, we allow the linear index [^3] We use a model with a linear index solely for the purpose of illustration. Our method applies to general models. $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ to take the following general form:
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\sum_{k=1}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
where $X_{k,it}$, $\theta_{k}$, $\alpha_{k,i}$, and $\gamma_{k,t}$ are the $k$-th components of $X_{it}$, $\theta$, $\alpha_{i}$, and $\gamma_{t}$, respectively, $k=1,\ldots,K$. In addition, when $X_{1,it}=1$, the DGP in (2.2) consists of the traditional individual- and time-effects in the intercept. When $X_{k,it}=Y_{it-1}$ for some $k$, the DGP is dynamic because each $Y_{it}$ depends on its past value.
Moreover, we also allow a subset of the coefficients to be homogeneous, thereby accommodating both heterogeneous and homogeneous parameters. For instance, when it is known that $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$, the model can be estimated under constraints $\alpha_{1,i}=\gamma_{1,t}=0$ for all $(i,t)$. In this case, the linear index is effectively
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=\theta_{1}X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
where $X_{1,it}$ is associated with a “homogeneous parameter” $\theta_{1}$ while all other $X_{k,it}$ ($k=2,\ldots,K$) are associated with heterogeneous parameters $(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})$. In a similar fashion, our approach also applies to, e.g.,
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})=(\alpha_{1,i}+\gamma_{1,t})X_{1,it}+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$
with the constraint $\theta_{1}=0$. Other variants of $\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ include:
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\alpha_{1,i}+\gamma_{1,t})+\theta_{2}Y_{it-1}+\sum_{k=3}^{K}\theta_{k}X_{k,it},$ (2.3)
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+\sum_{k=2}^{K}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it},$ (2.4)
$\pi(X_{it},\theta,\alpha_{i},\gamma_{t})$ $=(\theta_{1}+\alpha_{1,i}+\gamma_{1,t})+(\theta_{2}+\alpha_{2,i}+\gamma_{2,t})Y_{it-1}$
$\quad+\sum_{k=3}^{K_{1}}(\theta_{k}+\alpha_{k,i}+\gamma_{k,t})X_{k,it}+\sum_{k=K_{1}+1}^{K}\theta_{k}X_{k,it}$ (2.5)
for some $3\leq K_{1}<K$. To our knowledge, except (2.3) considered by Fernández-Val and Weidner (2016), none of the above specifications are fully addressed under nonlinear models with predetermined regressors. The specification in (2.4) is the same to Keane and Neal (2020) and Lu and Su (2023), except that they consider only the linear regression model. Jochmans and Otsu (2019) provide a bias correction technique for general nonlinear models with (2.4) only for strictly exogenous $X_{k,it}$. Currently, no methods are available for nonlinear models with (2.5).
We are interested in the estimation of and the inference about $\theta_{0}$ in (2.1), under the presence of the nuisance parameter $\phi^{0}$. Our approach is likelihood-based. Denote $\phi:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$ and define the log-likelihood function (“likelihood” hereafter) as
$\ell(\theta,\phi):=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}l_{i,t}(\theta,\phi),\qquad l_{i,t}(\theta,\phi):=\log f(Y_{it}|X_{it},\theta+\alpha_{i}+\gamma_{t}).$ (2.6)
Due to the additive structure $\theta+\alpha_{i}+\gamma_{t}$, $\theta$ is not identified because the likelihood function is invariant to transformations
$\alpha_{i}\mapsto\alpha_{i}+c_{1},\qquad\gamma_{t}\mapsto\gamma_{t}+c_{2},\qquad\theta\mapsto\theta-c_{1}-c_{2}$ (2.7)
for all pairs of constant $(c_{1},c_{2})$ and all $(i,t)$. To resolve this, we follow Lu and Su (2023) to reparameterize the likelihood by setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, thereby imposing the normalization $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$. Any “average effect” is absorbed by $\theta$. Let $\psi:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime},\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, the reparameterized likelihood can be constructed from the original $\ell\left(\theta,\phi\right)$ as
$\ell\left(\theta,D^{\prime}\psi\right)=:l(\theta,\psi),\qquad D:=\operatorname*{diag}\{D_{1},D_{2}\},$
where $D_{1}:=(\mathbb{I}_{N-1},-\iota_{N-1})\otimes\mathbb{I}_{K}$ and $D_{2}:=(\mathbb{I}_{T-1},-\iota_{T-1})\otimes\mathbb{I}_{K}$. This rules out the indeterminancy in (2.7) without explicitly using the Lagrangian. We use the following example to illustrate the reparameterization.
Example 2 (reparameterization)
For some scalar structural parameter $\mu_{0}$ of interests and $K=1$, consider $Y_{it}=\mu_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}$ where $\varepsilon_{it}$ is standard normal independent across both $i$ and $t$, and independent of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. We restrict the variance of $\varepsilon_{it}$ to be $1$ because we use this example only to demonstrate the identification of $\mu_{0}$. In Example 3 below, we indeed allow $\varepsilon_{it}$ to have an unknown variance.
Conditioning on $\phi^{0}$, the original likelihood function is
$\ell(\mu,\phi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{Nt}-\mu-\alpha_{N}-\gamma_{t})^{2}$
$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}(Y_{iT}-\mu-\alpha_{i}-\gamma_{T})^{2}-\frac{1}{NT}\frac{1}{2}(Y_{NT}-\mu-\alpha_{N}-\gamma_{T})^{2}.$
Setting $\alpha_{N}=-\sum_{i=1}^{N-1}\alpha_{i}$ and $\gamma_{T}=-\sum_{t=1}^{T-1}\gamma_{t}$, the reparameterized likelihood is
$l(\mu,\psi)$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{NT}\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}\frac{1}{2}(Y_{it}-\mu-\alpha_{i}-\gamma_{t})^{2}-\frac{1}{NT}\sum_{t=1}^{T-1}\frac{1}{2}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}-\gamma_{t}\right)^{2}$
$\quad-\frac{1}{NT}\sum_{i=1}^{N-1}\frac{1}{2}\left(Y_{iT}-\mu-\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}-\frac{1}{NT}\frac{1}{2}\left(Y_{NT}-\mu+\sum_{i=1}^{N-1}\alpha_{i}+\sum_{t=1}^{T-1}\gamma_{t}\right)^{2}.$
Calculating the first derivative of $l(\mu,\psi)$ w.r.t. $\alpha_{i}$, we have
$\partial_{\alpha_{i}}l(\mu,\psi)=\frac{1}{NT}\sum_{t=1}^{T}(Y_{it}-\mu-\alpha_{i})-\frac{1}{NT}\sum_{t=1}^{T}\left(Y_{Nt}-\mu+\sum_{i=1}^{N-1}\alpha_{i}\right)$
for every $i=1,\ldots,N-1$. Equating $\partial_{\alpha_{i}}l(\mu,\psi)=0$, we obtain
$\alpha_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\sum_{i=1}^{N-1}\alpha_{i}.$ (2.8)
Notice that (2.8) is independent of $\mu$ and all $\gamma_{t}$. Now, sum (2.8) over $i=1,\ldots,N-1$, we obtain
$\sum_{i=1}^{N-1}\alpha_{i}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}-\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}.$ (2.9)
Combining (2.8) and (2.9), it follows that
$\widehat{\alpha}_{i}=\frac{1}{T}\sum_{t=1}^{T}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad i=1,\ldots,N-1.$
It can be calculated in the same way that
$\widehat{\gamma}_{t}=\frac{1}{N}\sum_{i=1}^{N}Y_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad t=1,\ldots,T-1.$
Subsequently,
$\widehat{\mu}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$ (2.10)
In this simple example, $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ do not depend on $\mu$. This is generally not the case for complex models.
In addition, maximizing $l(\mu,\psi)$ is identical to maximizing the original likelihood $\ell\left(\mu,\phi\right)$ under constraints $\sum_{i=1}^{N}\alpha_{i}=0$ and $\sum_{t=1}^{T}\gamma_{t}=0$. Specifically, it can be verified that the Lagrangian $\ell\left(\mu,\phi\right)+\lambda^{\alpha}\sum_{i=1}^{N}\alpha_{i}+\lambda^{\gamma}\sum_{t=1}^{T}\gamma_{t}$, where $\lambda^{\alpha}$ and $\lambda^{\gamma}$ are the two Lagrange multipliers, delivers upon maximization the same $\widehat{\alpha}_{i}$ and $\widehat{\gamma}_{t}$ for $i=1,\ldots,N-1$ and $t=1,\ldots,T$, respectively. For $\widehat{\alpha}_{N}$ and $\widehat{\gamma}_{T}$, the Lagrangian produces
$\widehat{\alpha}_{N}=\frac{1}{T}\sum_{t=1}^{T}Y_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it},\qquad\widehat{\gamma}_{T}=\frac{1}{N}\sum_{i=1}^{N}Y_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}Y_{it}.$
Indeed, $\widehat{\alpha}_{N}=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i}$ as in Equation (2.9), and it can be verified that $\widehat{\gamma}_{T}=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t}$ as well.
2.2 Incidental Parameter Problem
The dynamic nonlinear heterogeneous parameter, DN-HP, models generally suffer from the incidental parameter problem (IPP). We now discuss this problem. We use $\mathbb{E}$ to denote the expectation w.r.t.
$\prod_{i=1}^{N}\prod_{t=1}^{T}f(y_{it}|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$
which is the true density evaluated at $\theta_{0}$, given: i) initial values of the predetermined regressors, ii) all strictly exogenous regressors, and iii) the unobserved effect $\phi^{0}$.
Remark 1 (expectation)
Our definition of $\mathbb{E}$ depends on the model specification and is the same to Arellano and Bonhomme (2009) and Fernández-Val and Weidner (2016). Consider the frequently used autoregressive model with $X_{it}=(Y_{it-1},Z_{it}^{\prime})^{\prime}$ for some strictly exogenous regressor $Z_{it}$. It is common that, for each $i=1,\ldots,N$, the density of $Y_{i1}$ is missing and (2.6) becomes a conditional likelihood as in Chapter 5 of Hamilton (1994). In this case, the expectation is taken w.r.t. $f_{Y_{i2},\ldots,Y_{iT}}\left(\cdot|Y_{i1},Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the density of only $Y_{i2},\ldots,Y_{iT}$, conditional on the initial value $Y_{i1}$, strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$, and the unobserved effects $\phi^{0}$. However, when the density of $Y_{i1}$ is indeed specified, so that (2.6) is an exact likelihood, the expectation is taken w.r.t. $f_{Y_{i1},\ldots,Y_{iT}}\left(\cdot|Z_{i1},\ldots,Z_{iT},\phi^{0};\theta_{0}\right)$, which is the joint density of all $Y_{i1},\ldots,Y_{iT}$, conditioning on the strictly exogenous regressors $Z_{i1},\ldots,Z_{iT}$ and the unobserved effects $\phi^{0}$ only.
Following the profiled likelihood framework of Pace and Salvan (2006), we continue our discussion by defining, for a given $\theta$ and each pair of $(N,T)$,
$\widehat{\psi}(\theta):=\arg\max_{\psi}l(\theta,\psi),\qquad\psi(\theta):=\arg\max_{\psi}\mathbb{E}[l(\theta,\psi)].$ (2.11)
Here $\psi(\theta)$ is referred to as the pseudo-true value and $l(\theta,\psi(\theta))$ is not subject to the IPP. See, e.g., Hahn and Newey (2004) and Arellano and Hahn (2016). In practice, however, $\psi(\theta)$ is generally infeasible and the plug-in version $l(\theta,\widehat{\psi}(\theta))$ is used instead. Specifically, the MLE of $\theta_{0}$, $\widehat{\theta}:=\arg\max_{\theta}\ell(\theta,\widehat{\psi}(\theta))$. The LR and LM test statistics (for hypothesis about $\theta_{0}$) are also constructed using $l(\theta,\widehat{\psi}(\theta))$. A fundamental problem is that $l(\theta,\widehat{\psi}(\theta))$ contains a non-negligible asymptotic bias of order $O(1/\sqrt{NT})$, due to the estimation errors in $\widehat{\psi}(\theta)$. This is referred to as the IPP in the context of the DN-HP model. The consequence is that, $\widehat{\theta}$ and many likelihood-based test statistics deviate from their standard asymptotic distribution, leading to incorrect inferences.
To demonstrate the non-negligible bias in $l(\theta,\widehat{\psi}(\theta))$, consider a Taylor expansion of $l(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$:
$l(\theta,\widehat{\psi}(\theta))=l(\theta,\psi(\theta))+s^{\prime}(\theta)[\widehat{\psi}(\theta)-\psi(\theta)]+\frac{1}{2}[\widehat{\psi}(\theta)-\psi(\theta)]^{\prime}H(\theta)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{l}(\theta),$ (2.12)
where $\widehat{\psi}(\theta)-\psi(\theta)$ represents the estimation errors of the fixed effects,
$s(\theta)$ $:=s(\theta,\psi(\theta)),\qquad s(\theta,\psi):=\partial_{\psi}l(\theta,\psi),$
$H\left(\theta\right)$ $:=H(\theta,\psi(\theta)),\qquad H(\theta,\psi):=\partial_{\psi\psi^{\prime}}l(\theta,\psi)$
are the score and Hessian. The remainder term $r^{l}(\theta)$ is characterized by Equation (A.17) in Appendix A. Here the estimation errors $\widehat{\psi}(\theta)-\psi(\theta)$ can be derived by another Taylor expansion. Seeing $s(\theta,\widehat{\psi}(\theta))=0$ as $\widehat{\psi}(\theta)$ is the maximizer, a Taylor expansion of $s(\theta,\widehat{\psi}(\theta))$ around $\psi(\theta)$ gives
00 0 $=s\left(\theta\right)+H\left(\theta\right)[\widehat{\psi}(\theta)-\psi(\theta)]+r^{s}(\theta),$
$\widehat{\psi}(\theta)-\psi(\theta)$ $=-[H\left(\theta\right)]^{-1}s\left(\theta\right)-[H\left(\theta\right)]^{-1}r^{s}(\theta),$ (2.13)
where $r^{s}(\theta)$ is the remainder term of the expansion, which is given by Equation (A.6) in Appendix A. Combining (2.13) and (2.12), and taking expectation, we obtain
$\sqrt{NT}\mathbb{E}[l(\theta,\widehat{\psi}(\theta))-l(\theta,\psi(\theta))]=-\sqrt{NT}B(\theta)+\sqrt{NT}\mathbb{E}r(\theta),\qquad B(\theta):=\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\},$ (2.14)
where $r(\theta)$ is a remainder term depending on $r^{l}(\theta)$ and $r^{s}(\theta)$, given by Equation (A.16) in Appendix A. By Lemma A.3 in Appendix A, we show that $\mathbb{E}r(\theta)$ is of order $o(1/\sqrt{NT})$, uniformly in $\theta$, as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$. This implies that $\sqrt{NT}\mathbb{E}r(\theta)=o(1)$ and is negligible. $B(\theta)$ is the IPP bias arising from the estimation errors of $\widehat{\psi}(\theta)$. As opposite to $\mathbb{E}r(\theta)$, the bias $B(\theta)$ is not negligible, in the sense that $\sqrt{NT}\mathbb{E}B(\theta)\not\rightarrow 0$ under the above asymptotic sequences $\left\{N,T\right\}$ (see Theorem 1). Consequently, $\widehat{\theta}$ and many likelihood-based test statistics inherit the bias from the likelihood, leading to invalid inferences. We consider the asymptotic sequences $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$, because $\sqrt{NT}B(\theta)$ explodes under $N/T\rightarrow 0$ or $N/T\rightarrow\infty$. We illustrate the IPP and its consequence to the structural parameter estimator in the example below, together with the issue of $N/T\rightarrow\kappa$.
Example 3 (heterogeneous mean model)
Following the same setting in Example 2 but let $\varepsilon_{it}$ be normal with mean 00 and variance $\sigma_{0}>0$. In the current example, $\theta_{0}=(\mu_{0},\sigma_{0})^{\prime}$ is the structural parameter of interests. This example is a modified version of Neyman and Scott (1948) and is similar to Fernández-Val and Weidner (2016), except that the mean of $Y_{it}$ is $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$, conditional on $\phi^{0}$.
Define [^4] In this simple example, $\varepsilon_{it}$ is independent across both $i$ and $t$. Therefore, $\mathbb{E}$ is simply w.r.t. the normal density having mean $\mu^{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ and variance $\sigma^{0}$. $\widetilde{Y}_{it}:=Y_{it}-\mathbb{E}Y_{it}$. Following the same steps in Example 2, we can derive
$\widehat{\alpha}_{i}-\alpha_{i}^{0}$ $=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad i=1,\ldots,N-1,$
$\widehat{\gamma}_{t}-\gamma_{t}^{0}$ $=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{it}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad t=1,\ldots,T-1,$ (2.15)
where $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ can be obtained by maximizing $\mathbb{E}l(\mu,\psi)$ in the same way. Denote additionally
$\widehat{\alpha}_{N}:=-\sum_{i=1}^{N-1}\widehat{\alpha}_{i},\qquad\alpha_{N}^{0}:=-\sum_{i=1}^{N-1}\alpha_{i}^{0},\qquad\widehat{\gamma}_{T}:=-\sum_{t=1}^{T-1}\widehat{\gamma}_{t},\qquad\gamma_{T}^{0}:=-\sum_{t=1}^{T-1}\gamma_{t}^{0}.$
It follows that
$\widehat{\alpha}_{N}-\alpha_{N}^{0}=\frac{1}{T}\sum_{t=1}^{T}\widetilde{Y}_{Nt}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it},\qquad\widehat{\gamma}_{T}-\gamma_{T}^{0}=\frac{1}{N}\sum_{i=1}^{N}\widetilde{Y}_{iT}-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$ (2.16)
Now, for $\psi^{0}:=({\alpha}_{1}^{0},\ldots,{\alpha}_{N-1}^{0},{\gamma}_{1}^{0},\ldots,{\gamma}_{T-1}^{0})^{\prime}$, $\widehat{\psi}:=(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{N-1},\widehat{\gamma}_{1},\ldots,\widehat{\gamma}_{T-1})^{\prime}$, and $Y_{it}(\mu):=Y_{it}-\mu-\alpha_{i}^{0}-\gamma_{t}^{0}$, we can calculate
$l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\alpha}_{i}-\alpha_{i}^{0})\right)(\widehat{\alpha}_{i}-\alpha_{i}^{0})$
$\quad\quad+\frac{1}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(Y_{it}(\mu)-\frac{1}{2}(\widehat{\gamma}_{t}-\gamma_{t}^{0})\right)(\widehat{\gamma}_{t}-\gamma_{t}^{0}),$ (2.17)
where $\theta=(\mu,\sigma)^{\prime}$ with $\sigma>0$. Substituting (2.15) and (2.16) into (2.17), we obtain, after some algebra,
$l(\theta,\widehat{\psi})-l(\theta,\psi^{0})$ $=\frac{1}{2\sigma}\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}+\frac{1}{2\sigma}\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
$\quad\quad-\frac{2}{\sigma}\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\left(\mu-\mu^{0}\right)\widetilde{Y}_{it}+O_{p}(1/(NT)).$
Here, $\mathbb{E}\widetilde{Y}_{it}=0$ by definition and $\mathbb{E}\widetilde{Y}_{it}\widetilde{Y}_{js}=0$, for $i\neq j$ or $t\neq s$, by independence. Therefore, calculating the expectation, we arrive at
$\mathbb{E[}l(\theta,\widehat{\psi})-l(\theta,\psi^{0})]=-B(\theta)+O(1/(NT)),\qquad B(\theta)=-\left(\frac{1}{T}+\frac{1}{N}\right)\frac{\sigma^{0}}{2\sigma}.$
The term $B(\theta)$ is the IPP bias. Apparently,
$\sqrt{NT}B(\theta)=-\left(\sqrt{\frac{N}{T}}+\sqrt{\frac{T}{N}}\right)\frac{\sigma^{0}}{2\sigma}\not\rightarrow 0$
as $N,T\rightarrow\infty$ and $N/T\rightarrow\kappa$. When $N/T\rightarrow 0$ (or $N/T\rightarrow\infty$), $\sqrt{T/N}\rightarrow\infty$ (or $\sqrt{N/T}\rightarrow\infty$) and $\sqrt{NT}B(\theta)$ explodes.
The bias term $B(\theta)$ causes the variance estimator to be biased. We illustrate this consequence here. Observe that
$l(\theta,\widehat{\psi})$ $=-\frac{1}{2}\log(2\pi)-\frac{1}{2}\log\sigma-\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2},$
$\partial_{\sigma}l(\theta,\widehat{\psi})$ $=-\frac{1}{2\sigma}+\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{1}{2\sigma^{2}}(Y_{it}-\mu-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}.$
Following the same steps in Example 2, we obtain exactly the same estimator of $\mu_{0}$, $\widehat{\mu}$, as in Equation (2.10); and it is independent of $\sigma$. Therefore, equating $\partial_{\sigma}l(\theta,\widehat{\psi})=0$, substituting $\mu$ with $\widehat{\mu}$, and solving for $\sigma$, the estimator of $\sigma^{0}$ is
$\widehat{\sigma}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}(Y_{it}-\widehat{\mu}-\widehat{\alpha}_{i}-\widehat{\gamma}_{t})^{2}$
$=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}[Y_{it}-\mu_{0}-\alpha_{i}^{0}-\gamma_{t}^{0}-(\widehat{\mu}-\mu_{0})-(\widehat{\alpha}_{i}-\alpha_{i}^{0})-(\widehat{\gamma}_{t}-\gamma_{t}^{0})]^{2},$
where it is not difficult to see
$\widehat{\mu}-\mu_{0}=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}.$
Therefore, seeing $\widetilde{Y}_{it}=Y_{it}-\mu_{0}-\alpha_{i}^{0}-\gamma_{t}^{0}$ by definition, we finally obtain
$\widehat{\sigma}$ $=\frac{1}{NT}\sum_{i=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}^{2}-\frac{1}{NT^{2}}\sum_{i=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{is}-\frac{1}{N^{2}T}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\widetilde{Y}_{it}\widetilde{Y}_{jt}$
$\quad\quad+\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{t=1}^{T}\sum_{s=1}^{T},\widetilde{Y}_{it}\widetilde{Y}_{js}$
from where we conclude that
$\mathbb{E}\widehat{\sigma}=\sigma_{0}-\frac{\sigma_{0}}{T}-\frac{\sigma_{0}}{N}+\frac{\sigma_{0}}{NT},$
leading to
$\sqrt{NT}(\mathbb{E}\widehat{\sigma}-\sigma_{0})\rightarrow-\kappa\sigma_{0}-\frac{\sigma_{0}}{\kappa}$
as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$. While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $-\sigma_{0}/T$ and $-\sigma_{0}/N$. These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $-\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
3 Bias Correction and Asymptotic Theory
The bias expansion of $l(\theta,\widehat{\psi}(\theta))$ in Equation (2.14) and Example 3 above motivate our approach of correcting the likelihood $l(\theta,\widehat{\psi}(\theta))$ by adding $B(\theta)$. The corrected likelihood can be viewed as an approximation to the IPP-free infeasible likelihood $l(\theta,\psi(\theta))$, with approximation error $o(1/\sqrt{NT})$. In Theorem 1 below we formalize the expansion in Equation (2.14). Specifically, we show that $B(\theta)$ can be further decomposed into $B(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT})$, in which $B_{\alpha}(\theta)$ appears because of the estimation errors of the individual-specific heterogeneities $\alpha_{i}^{0}$, while $B_{\gamma}(\theta)$ is present due to the estimation errors of the time-specific heterogeneities $\gamma_{t}^{0}$. The exact formulas for both terms are given below. Theorem 2 follows to establish the corrected likelihood as
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)$ (3.1)
where we denote $\widehat{l}(\theta):=l(\theta,\widehat{\psi}(\theta))$ for simplicity from now on. $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are estimates of $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, respectively. The analytical formulas for $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are provided below. We then show that
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
where we use $l(\theta):=l(\theta,\psi(\theta))$ for simplicity below. This implies that $L(\theta)$ does not suffer from the IPP. The corrected estimator of $\theta_{0}$, and the LR, LM, and Wald test statistics, may subsequently be derived from $L\left(\theta\right)$, instead of $\widehat{l}(\theta)$. Specifically, the corrected estimator of $\theta_{0}$ is
$\widehat{\theta}_{L}:=\arg\max_{\theta}L(\theta).$
Under mild regularity conditions, $\widehat{\theta}_{L}$ is asymptotically unbiased, and its asymptotic properties align with the standard results of the classical MLE without the nuisance parameters. Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis. These results are formally stated in Corollaries 1 and 2, respectively.
Remark 2 (computation)
The computation of $\widehat{\theta}_{L}$ involves a two-layer maximization but can be done efficiently. In the inner layer, $\widehat{\psi}(\theta)$ is obtained by maximizing $l(\theta,\psi)$, keeping $\theta$ fixed. This is large-scale maximization problem and a few strategies can be exploited to expedite the procedure:
  1. 1.
    For many commonly used models, the first and second derivatives of $l(\theta,\psi)$ have clear analytical expressions, so that the standard Newton-Raphson algorithm can be employed to obtain $\widehat{\psi}(\theta)$.
  2. 2.
    Many models (specifically, generalized linear models) are equipped with linear indices, as in Example 1. In that case, the chain-rule-based technique in Sun and Dhaene (2019) can be used to achieve a fast calculation of the derivatives. The computational complexity of such a technique only grows with the number of linear indices, not the dimension of $\psi$, even without analytical expressions for the derivatives of $l(\theta,\psi)$.
  3. 3.
    The derivatives of the reparameterized $l(\theta,\psi)$ can be identically constructed from the derivatives of the original likelihood $\ell(\theta,\phi)$. Specifically, the reparameterization implies
    $\partial_{\psi}l(\theta,\psi)=D\partial_{\phi}\ell(\theta,\phi),\qquad\partial_{\psi\psi^{\prime}}l(\theta,\psi)=D\partial_{\phi\phi^{\prime}}\ell(\theta,\phi)D^{\prime},$
    holding true for every $\phi=D^{\prime}\psi$ and every $\theta$.
In the outer layer where $L(\theta)$ is maximized w.r.t. $\theta$, it can be tedious to obtain analytical formulas for the derivatives of the corrected likelihood $L(\theta)$. However, the dimension of $\theta$ is usually small, so that invoking the Newton-Raphson algorithm with numerical derivatives generally does not suffer from too much computational pressure, provided that the inner-layer maximization is done efficiently.
Now, we proceed to our bias correction procedure on the likelihood and present the related asymptotic theory. Denote $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta)$ and $\phi(\theta):=D^{\prime}\psi(\theta)$, which are essentially the “unparameterized” counterparts of $\widehat{\psi}(\theta)$ and $\psi(\theta)$, respectively. We impose the following assumption.
Assumption 1
  • (i) Suppose $N/T\rightarrow\kappa$ for some $0<\kappa<\infty$ as $N,T\rightarrow\infty$.
[…]
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Phi}\left|\partial_{\phi_{1}\cdots\phi_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $\phi_{s}$ represents any element of $\phi$ and $S\in\{0,1,2,3\}$ ($S=0$ is understood as not taking derivatives), and
    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\{[g(w_{it})]^{\eta}\}<\infty,$
    for some $\eta>2$.
  • (iii) For each $k=1,\ldots,K$, $\sum_{i=1}^{N}\alpha_{k,i}^{0}=\sum_{t=1}^{T}\gamma_{k,t}^{0}=0$, where $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ are the $k$-th component of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$, respectively.
  • (iv)  For each $\theta\in\Theta$, $\mathbb{P}[l(\theta,\psi)\neq l(\theta,\psi(\theta))]>0$ for every $\psi$ such that $D^{\prime}\psi\in\Phi$ and $\psi\neq\psi(\theta)$.
  • (v) Conditional on $\phi^{0}$, $\{w_{it}\}$ are independent across $i$ and conditionally strong mixing with mixing coefficient
    $a_{i}(m):=\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}(A\cap B|\phi^{0})-\mathbb{P}(A|\phi^{0})\mathbb{P}(B|\phi^{0})\right|$
    such that, for $\eta>2$ as above,
    $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}[a_{i}(m)]^{1-2/\eta}<\infty,$
    where $\mathcal{A}_{it}$ and $\mathcal{B}_{it}$ are the $\sigma$-algebras generated by $\{w_{is}:1\leq s\leq t\}$ and $\{w_{is}:t\leq s\leq T\}$, respectively.
  • (vi) The matrix has $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
    $\sup_{N>N_{0},T>T_{0}}\max_{1\leq p\leq(N+T-2)K}\sup_{\theta\in\Theta}\lambda_{p}(\theta)<0$
    for some integers $N_{0}$ and $T_{0}$.
Here Assumption 1 is the same to Fernández-Val and Weidner (2016). It is used to ensure that the bias $\sqrt{NT}B(\theta)$ does not explode as $N,T\rightarrow\infty$, as discussed above. Assumption 1 is standard in the literature. See, e.g., Hahn and Newey (2004), Hahn and Kuersteiner (2011), and Arellano and Hahn (2016). It simply requires that the likelihood function $l_{i,t}(\theta,\phi)$ is differentiable in $\phi$ at least three times. The term $B(\theta)$ in (2.14) depends on the first and second derivatives, and the remainder $r(\theta)$ depends on the third. We require $l_{i,t}(\theta,\phi)$ and its derivatives to be uniformly bounded by a function $g(w_{it})$ whose moment exists to an order greater than the second. This is used to ensure, amongst other things, i) the interchangeability of $\mathbb{E}$ and $\partial$; and ii) the existence of moments, such as $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ in (2.14), for every $\theta\in\Theta$. It is also used to establish the uniform rate of convergence, over $\theta\in\Theta$, of the corrected likelihood $L(\theta)$. The uniform convergence of $L(\theta)$ guarantees that the bias correction effect carries over to both the estimators and likelihood-based test statistics. Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 00 over $i=1,\ldots,N$ and $t=1,\ldots,T$. Given the additive structure of $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ in (2.1), this condition is without loss of generality because we merely let $\theta_{0}$ capture the average effects of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Assumption 1 is an identification condition for $\psi(\theta)$. It is standard for maximum likelihood models. See, e.g., Newey and McFadden (1994). Assumption 1 is common in the field of dynamic panels. See, for instance, Hahn and Kuersteiner (2007), Fernández-Val and Weidner (2016), and Dhaene and Jochmans (2015). Notice that, the argument is conditional on the true heterogeneous effects $\phi^{0}$. Without conditioning, $w_{it}$ may not be, for instance, independent over $i$, because $w_{it}$ and $w_{jt}$ may depend on the same $\gamma_{t}^{0}$. In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $[a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$. The mixing condition is used to guarantee that certain autocovariances which appear in our proof are summable by Davydov’s inequality. Notice that we are not requiring $w_{it}$ to be identically distributed over $i$ or stationary over $t$. These are generally not possible because $w_{it}$ depends on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 00 for all $N>N_{0}$ and $T>T_{0}$. This ensures that $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ is well defined for large $N$ and $T$. This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood. We state the condition with “less than 00” because the Hessian in a maximum likelihood model is usually negative definite.
Next, partition [^5] Note that $s(\theta)=s(\theta)-\mathbb{E}s\left(\theta\right)$ because $\mathbb{E}s\left(\theta\right)=0$ by definition. We keep $\mathbb{E}s\left(\theta\right)$ here for the convenience of constructing the corrected likelihood. $s(\theta)-\mathbb{E}s\left(\theta\right)$ and $[\mathbb{E}H(\theta)]^{-1}$ as
$s(\theta)-\mathbb{E}s\left(\theta\right)=:\left(\begin{array}[c]{c}\widetilde{s}_{\alpha}(\theta)\\ \widetilde{s}_{\gamma}(\theta)\end{array}\right),\qquad[\mathbb{E}H(\theta)]^{-1}=:\left(\begin{array}[c]{cc}H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right),$ (3.2)
where $\widetilde{s}_{\alpha}(\theta)$ is $(N-1)K\times 1$ and $\widetilde{s}_{\gamma}(\theta)$ is $(T-1)K\times 1$ while $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ is $(N-1)K\times(N-1)K$ and $H_{\gamma\gamma}^{\ast}(\theta)$ is $(T-1)K\times(T-1)K$. The off-diagonal blocks $H_{\alpha\gamma}^{\ast}(\theta)$ and $H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)$ are defined accordingly. Further denote
$S_{\alpha\alpha}(\theta):=\mathbb{E\{}\widetilde{s}_{\alpha}(\theta)\mathbb{[}\widetilde{s}_{\alpha}(\theta)]^{\prime}\},\qquad S_{\gamma\gamma}(\theta):=\mathbb{E\{}\widetilde{s}_{\gamma}(\theta)\mathbb{[}\widetilde{s}_{\gamma}(\theta)]^{\prime}\}.$
The next theorem formally states the asymptotic expansion of $\widehat{l}(\theta)$.
Theorem 1
Under Assumption 1 and uniformly over $\theta\in\Theta$,
$\mathbb{E}l(\theta)=\mathbb{E}\widehat{l}(\theta)+B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT}),$
as $N,T\rightarrow\infty$, where
$B_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
[…]
The bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ involves $\psi(\theta)$ and $\mathbb{E}$ which are infeasible to compute. However, $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ can be estimated by plug-in estimates of $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $H_{\gamma\gamma}^{\ast}(\theta)$, $S_{\alpha\alpha}(\theta)$, and $S_{\gamma\gamma}(\theta)$. The two Hessian terms can be estimated in a straightforward way. Denote $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta))$ and partition $[\widehat{H}(\theta)]^{-1}$ into
$[\widehat{H}(\theta)]^{-1}=:\left(\begin{array}[c]{cc}\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ \widehat{H}_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right)$ (3.3)
in the same manner as in (3.2). The term $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ can be estimated by the $(N-1)K\times(N-1)K$ block $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, and $H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$ by the $(T-1)K\times(T-1)K$ block $\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$. Estimators of $S_{\alpha\alpha}(\theta)$ and $S_{\gamma\gamma}(\theta)$ are constructed from the derivatives of $l_{i,t}(\theta,\phi)$, the original likelihood for individual $i$ at time $t$. This procedure needs some explanation. The reparameterization ensures the following identity:
$S_{\alpha\alpha}(\theta)=D_{1}\mathcal{S}_{\alpha\alpha}(\theta)D_{1}^{\prime},$
where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)=D^{\prime}\psi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t. $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$. Specifically,
$\mathcal{S}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E\{}\widetilde{s}_{t}^{\alpha}(\theta)[\widetilde{s}_{s}^{\alpha}(\theta)]^{\prime}\},\qquad\widetilde{s}_{t}^{\alpha}(\theta):=\{[\widetilde{s}_{1,t}^{\alpha}(\theta)]^{\prime},\ldots,[\widetilde{s}_{N,t}^{\alpha}(\theta)]^{\prime}\}^{\prime},$
$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)),\qquad s_{i,t}^{\alpha}(\theta,\phi):=\partial_{\alpha_{i}}l_{i,t}(\theta,\phi).$
Here, $\widetilde{s}_{i,t}^{\alpha}(\theta)$ is the centered version of $s_{i,t}^{\alpha}(\theta,\phi(\theta))$, the derivative, w.r.t. $\alpha_{i}$ and evaluated at $\phi(\theta)$, of the original likelihood $l_{i,t}(\theta,\phi)$ for individual $i$ at time $t$. Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
$\mathcal{S}_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\operatorname*{diag}\{\widetilde{s}_{1,t}^{\alpha}(\theta)[\widetilde{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widetilde{s}_{N,t}^{\alpha}(\theta)[\widetilde{s}_{N,s}^{\alpha}(\theta)]^{\prime}\}.$
Naturally, the estimator of $S_{\alpha\alpha}(\theta)$ is
$\widehat{S}_{\alpha\alpha}(\theta)$ $:=D_{1}\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)D_{1}^{\prime},$ (3.4)
$\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\{\left|t-s\right|\leq\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\},$
$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta)).$
The indicator $\mathds{1}\{\left|t-s\right|\leq\tau\}$ is a truncation mechanism with truncation parameter $\tau$ (see Remark 3 below). Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
$S_{\gamma\gamma}(\theta)=D_{2}\mathcal{S}_{\gamma\gamma}(\theta)D_{2}^{\prime},$
where $\mathcal{S}_{\gamma\gamma}(\theta)$ is the outer product of the $TK\times 1$ vector of $\partial_{\gamma}\ell(\theta,\phi)-\mathbb{E}\partial_{\gamma}\ell(\theta,\phi)$, for $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$, evaluated at $\phi=\phi(\theta)$:
$\mathcal{S}_{\gamma\gamma}(\theta)$ $=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\},\qquad\widetilde{s}_{i}^{\gamma}(\theta):=\{[\widetilde{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widetilde{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)),\qquad s_{i,t}^{\gamma}(\theta,\phi):=\partial_{\gamma_{t}}l_{i,t}(\theta,\phi),$
where $\widetilde{s}_{i,t}^{\gamma}(\theta)$ is the centered version of $s_{i,t}^{\gamma}(\theta,\phi(\theta))$, the derivative of $l_{i,t}(\theta,\phi)$ w.r.t. $\gamma_{t}$ and evaluated at $\phi(\theta)$. By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (3.5)
Intuitively, the estimator for $S_{\gamma\gamma}(\theta)$ is
$\widehat{S}_{\gamma\gamma}(\theta)$ $=D_{2}\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)D_{2}^{\prime},$ (3.6)
$\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}\widehat{s}_{i}^{\gamma}(\theta)[\widehat{s}_{i}^{\gamma}(\theta)]^{\prime}\},\qquad\widehat{s}_{i}^{\gamma}(\theta):=\{[\widehat{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widehat{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta)).$
Remark 3 (truncation)
In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016). This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case. Detailed discussions in this regard can be found in Andrews (1991). Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2). Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2 below). Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$. For intermediate $T$, we recommend following the advices from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values. In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (3.5) using cross-sectional independence. One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 00. We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics. The comparison is, hence, not reported in the paper.
Using Equations (3.3), (3.4), and (3.6), we can construct the corrected likelihood $L(\theta)$, which is stated in the theorem below with its rate of convergence.
Theorem 2
Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta),$
as in (3.1), with
$\widehat{B}_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[\widehat{S}_{\alpha\alpha}(\theta)\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad\widehat{B}_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}\widehat{S}_{\gamma\gamma}(\theta)\widehat{H}_{\gamma\gamma}^{\ast}(\theta)].$
As mentioned above, Theorem 2 states that $L(\theta)$ does not suffer from the IPP. Showing 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1. Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error. For the truncation, we show that $1/\left(N^{2}T^{2}\right)\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\cdot\right)$, which is the truncated (i.e., the removed) part in (3.4), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions. Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
Next, we turn our attention to the asymptotic properties of the $L(\theta)$-based estimators $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta)$ and of the $L(\theta)$-based LR, LM, Wald test statistics. We enforce the following assumption.
Assumption 2
  • (i) The likelihood $l_{i,t}(\theta,\phi)$ is four-time continuously differentiable w.r.t. to both $\theta\in\Theta$ and $\phi\in\Phi$ almost surely, such that
    $\sup_{\theta\in\Theta}\sup_{\phi\in\Phi}\left|\partial_{\nu_{1}\cdots\nu_{S}}l_{i,t}(\theta,\phi)\right|<g(w_{it}),$
    where $g(w_{it})$ is defined in Assumption 1, $\nu_{s}$ represents any element of $(\theta^{\prime},\phi^{\prime})^{\prime}$, and $S\in\{1,2,3,4\}$.
  • (ii) For every $\theta\in\Theta$ and $\theta\neq\theta_{0}$, $\mathbb{P}[l(\theta)\neq l(\theta_{0})]>0$.
  • (iii) For large $N$ and $T$,
    $\mathbb{E}\sup_{\theta\in\Theta}\left|l(\theta)\right|<\infty.$
  • (iv) The sequence of functions $\{\left\|\nabla_{\theta}l(\theta)\right\|:\theta\in\Theta\}$ is uniformly integrable, and
    $\mathbb{E}\sup_{\theta\in\Theta}\left\|\nabla_{\theta\theta^{\prime}}l(\theta)\right\|<\infty$
[…]
  • (v) The profiled Hessian $\mathbb{E}\nabla_{\theta\theta^{\prime}}l(\theta)$ is strictly negative definite for large $N$ and $T$.
Here Assumption 2 is similar to Assumption 1, except that we impose that $l_{i,t}(\theta,\phi)$ is four-time differentiable, and require that the derivatives of $l_{i,t}(\theta,\phi)$ w.r.t. to both $\theta$ and $\phi$ to be uniformly dominated by $g(w_{it})$. This assumption ensures that the corrected likelihood $L(\theta)$, which depends on second derivatives of $l_{i,t}(\theta,\phi)$, is twice differentiable. This is required by standard techniques used to show, e.g., the asymptotic normality of $\widehat{\theta}_{L}$. See., e.g., Hayashi (2000). Similar to Assumption 1, Assumption 2 is an identification condition for $\theta_{0}$. It guarantees that $\operatorname*{plim}_{N,T\rightarrow\infty}l(\theta)$ is uniquely maximized at $\theta_{0}$. Assumption 2 is a dominance condition implying that $l(\theta)-\mathbb{E}l(\theta)$ converges to 00 uniformly as $N,T\rightarrow\infty$. Together, Assumptions 2 to 2 are used to establish the consistency of $\widehat{\theta}_{L}$. In addition to these, Assumption 2 establishes a dominance condition for the profiled Hessian, so that $\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})\overset{\mathbb{P}}{\longrightarrow}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]$ for any consistent estimator $\overline{\theta}$ of $\theta_{0}$. Assumption 2 requires that the expected profiled Hessian is strictly negative definite for sufficiently large $N$ and $T$. Assumptions 2 and 2 are additionally imposed to establish the asymptotic distributions for the quantities of interests.
The corollary below states the asymptotic properties of $\widehat{\theta}_{L}$.
Corollary 1
Under Assumptions 1 and 2, $\widehat{\theta}_{L}\overset{\mathbb{P}}{\rightarrow}\theta_{0}$ and
$\sqrt{NT}[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K})$
as $N,T\rightarrow\infty$.
Here, $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ is the square root of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ which can be obtained by square-rooting the eigenvalues in the eigendecomposition of $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$. By Assumption 2, $\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ is strictly negative definite so that $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ only has strictly positive eigenvalues. In addition, since we focus on maximum likelihood models, we use the information matrix equality to simplify the presentation of Corollary 1. In case the information matrix equality does not hold [^6] The information matrix equality generally does not hold for, e.g., non-likelihood M-estimation models, misspecified models, and quasi-likelihood models., $-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})$ shall be replaced by its sandwich form and $[-\mathbb{E}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]^{1/2}$ shall be constructed thereof. See, e.g., White (1982) for detailed discussions.
For inference procedures, we consider a generic null hypothesis $H_{0}:R(\theta_{0})=0$, where $R(\theta)$ is a known $r\times 1$ ($r\leq K$) vector-valued non-random function with Jacobian $J(\theta):=\partial_{\theta^{\prime}}R(\theta)$ satisfying $\operatorname*{rank}[J(\theta)]=r$. The corollary below gives the asymptotic distribution, under $H_{0}$, of $\widehat{\xi}_{LR}$ (the LR statistic), $\widehat{\xi}_{LM}$ (the LM statistic), and $\widehat{\xi}_{Wald}$ (the Wald statistic), defined respectively as
$\widehat{\xi}_{LR}$ $:=-2NT\{L(\widehat{\theta}_{R})-L(\widehat{\theta}_{L})\},$
$\widehat{\xi}_{LM}$ $:=-NT\{\triangledown_{\theta^{\prime}}L(\widehat{\theta}_{R})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\triangledown_{\theta}L(\widehat{\theta}_{R})\},$
$\widehat{\xi}_{Wald}$ $:=-NT\{R^{\prime}(\widehat{\theta}_{L})[J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J^{\prime}(\widehat{\theta}_{L})]^{-1}R(\widehat{\theta}_{L})\},$
where $\widehat{\theta}_{R}:=\arg\max_{\theta}L(\theta)$ subject to $R(\theta)=0$.
Corollary 2
Under Assumptions 1 to 2 and $H_{0}$,
$\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}(r)$
as $N,T\rightarrow\infty$, where $\chi^{2}(r)$ is the $\chi^{2}$-distribution with degrees of freedom $r$.
Similarly, the test statistics $\widehat{\xi}_{LM}$ and $\widehat{\xi}_{Wald}$ are constructed under the information matrix equality. Adjustments are generally required if it does not hold. For instance, White (1982) shows that $-[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}$ involved in $\widehat{\xi}_{Wald}$ shall be replaced by its sandwich form for quasi-likelihood models.
Given Theorem 2, Corollaries 1 and 2 are established using standard techniques. Particularly, Theorem 2 guarantees $L(\theta)\rightarrow_{p}\mathbb{E}l(\theta)$, so that the consistency of $\widehat{\theta}_{L}$ follows from Newey and McFadden (1994). Theorem 2 also implies that $\nabla_{\theta}L(\theta)-\nabla_{\theta}l(\theta)$ is $o_{\mathbb{P}}(1/\sqrt{NT})$, from which the asymptotic normality is established using techniques in, e.g., Hayashi (2000). The proof of Corollary 2 also closely resembles the standard techniques in, e.g., Hayashi (2000), where it is shown that all $\widehat{\xi}_{LR}$, $\widehat{\xi}_{LM}$, and $\widehat{\xi}_{Wald}$ are equivalent, converging in probability to some $\widehat{\xi}$ which is asymptotically $\chi^{2}(r)$.
4 Simulation
We conduct the simulation study in this section. We first illustrate that our procedure is capable of reducing the bias of the MLE as well as delivering likelihood-based tests with an empirical size very close to the nominal level. Followed by this, we then consider a different design in which the MLEs and the tests suffer from severe biases and size-distortions. We demonstrate that our approach leads to dramatic improvements on the estimators and substantial reductions of the Type-I error risks of the tests, even in a severely biased setting. Finally in this section, we compare our method with the Jackknife [^7] Since there has currently been no Jackknife procedures for DN-HP models. We compared our method to the Jackknife of Fernández-Val and Weidner (2016), who consider models with fixed effects only in the intercept. and show that our method out-performs.
In Design 1, we consider a dynamic binary response model as:
$Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ (4.1)
where the true value is $\theta_{0}=(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ and the disturbance $\varepsilon_{it}$ is standard logistic, independent and identically distributed (i.i.d.) over $(i,t)$ and independent from $X_{it}$. The true incidental parameter $\{\alpha_{1,i}^{0},\gamma_{1,t}^{0}\}$ are generated from $\mathcal{N}(0,0.01)$, and demeaned after they are generated. $\{\alpha_{2,i}^{0},\gamma_{2,t}^{0}\}$ are also generated from $\mathcal{N}(0,0.01)$ but only $\gamma_{2,t}^{0}$ is demeaned. Conditional on the generated $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ ($k=1,2$), the regressor $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$ i.i.d. over $(i,t)$ and the initial $Y_{i0}=\mathds{1}\{(\beta_{0}+\alpha_{1,i}^{0})X_{i0}+\alpha_{2,i}^{0}+\varepsilon_{i0}>0\}$ with $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$ i.i.d. over $(i,t)$. We also simulate a static model, where we set $\rho\equiv 0$ in (4.1), using data generated in the same way but setting $\rho_{0}=0$ (hence $\theta_{0}=\beta_{0}$).
Table 1 reports the estimation bias and RMSE, under Design 1, of $\widehat{\theta}_{L}$ and $\widehat{\theta}$ for both the static ($\theta_{0}=\beta_{0}$) and dynamic ($\theta_{0}=(\rho_{0},\beta_{0})^{\prime}$) logit models. For the static model, $\widehat{\beta}$ denotes the uncorrected MLE obtained by maximing $\widehat{l}(\theta)$, and $\widehat{\beta}_{L}$ is the corrected estimator, obtained by maximing $L(\theta)$. For the dynamic model, $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimator for $(\rho_{0},\beta_{0})^{\prime}$, while $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents the bias-corrected estimator, with the truncation parameter $\tau\in(1,2)$. For the static model (with $\rho_{0}\equiv 0$), we set $\tau=0$ since the data are independent over $i$, conditional on $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}$. The results reveal that the uncorrected estimators are severely biased in both models, while our bias-corrected estimators substantially reduce the bias. For instance, in the static model with $(N,T)=(30,30)$, the bias for $\widehat{\beta}_{L}$ is only $0.05$, which is less than $1/3$ of the significant bias of $0.158$ for $\widehat{\beta}$. As the sample size increases to $(N,T)=(90,90)$, $\widehat{\beta}_{L}$ is nearly unbiased, while the bias of $\widehat{\beta}$ remaining five times larger than that of $\widehat{\beta}_{L}$. Furthermore, the RMSEs from our corrected estimators are consistently lower than those of the uncorrected estimators, especially with smaller sample sizes. This indicates that our procedure is capable of reducing the IPP bias without increasing the RMSE. For large $(N,T)$, the choice of $\tau$ has negligible impact on the performance of our bias-corrected estimators.
Table 1: Comparisons of Uncorrected and Corrected Logit Estimates, Design 1
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.158$ $0.253$ $0.075$ $0.109$ $0.051$ $0.072$
$\widehat{\beta}_{L}$ $0.050$ $0.175$ $0.018$ $0.073$ $0.010$ $0.049$
Dynamic Model
$\widehat{\beta}$ $0.165$ $0.261$ $0.074$ $0.111$ $0.046$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.181$ $0.017$ $0.077$ $0.006$ $0.051$
$\widehat{\beta}_{L}^{(2)}$ $0.061$ $0.183$ $0.018$ $0.078$ $0.007$ $0.051$
$\widehat{\rho}$ $-0.107$ $0.201$ $-0.048$ $0.093$ $-0.033$ $0.058$
$\widehat{\rho}_{L}^{(1)}$ $-0.044$ $0.159$ $-0.014$ $0.076$ $-0.010$ $0.047$
$\widehat{\rho}_{L}^{(2)}$ $-0.038$ $0.161$ $-0.009$ $0.077$ $-0.006$ $0.046$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Next, under Design 1, we conduct LR tests using both the corrected and the uncorrected likelihoods for the null hypothesis $H_{0}:\theta_{0}=(0.5,0.5)^{\prime}+\delta$ with $\delta\in\{0,\pm 0.2,\pm 0.15,\pm 0.1\}$. We report the empirical size ($\delta=0$) and the empirical power ($\delta\neq 0$) at the $5\%$ significance level. For the corrected likelihood, we set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model. Table 2 presents the empirical size. The results show that the LR test using the uncorrected likelihood suffers from a severe size distortion. On the opposite, the empirical size from our bias-corrected likelihood maintains the significance level as the sample size increases.
Table 2: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 1
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $16\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$. The number of replications is $1000$.
Table 3 presents the empirical power of LR tests based on our corrected likelihood under Design 1. The empirical power increases as both the value of $\delta$ and the sample size grow. For $(N,T)=(90,90)$, when $\delta=\pm 0.2$, the rejection rates reach nearly $100\%$ for both the static and dynamic models. We also perform LM and Wald test, reporting the empirical size and power under the same setting in Appendix B.1. The results align with those of the LR test. We additionally perform the simulations for a probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$ i.i.d. over $(i,t)$ and independent from $X_{it}$. The results are provided in Appendix B.2.
Table 3: Empirical Power of LR Test for Logit Model using $L(\theta)$, Design 1
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $45\%$ $32\%$ $23\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $63\%$ $37\%$ $25\%$ $49\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $33\%$ $21\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $95\%$ $74\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Next, we conduct the same experiment under Design 2, where we generate $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$, keeping everything else intact as in (4.1). We increase the volatility of $\alpha_{k,i}^{0}$, $\gamma_{k,t}^{0}$, $X_{it}$, and $X_{i0}$ to induce larger biases and size distortions. This time, we present the results for the probit model and leave that for the logit model in Appendix B.1.
Table 4 presents the comaprison of the bias and RMSE of the corrected and the uncorrected estimators for the probit model, under Design 2. First of all, the probit estimator here is much more biased than the previous logit counterparts. For the static model with $(N,T)=(30,30)$, $\widehat{\beta}$ contains a bias of $0.943$. Recall that the true value is only $\beta_{0}=0.5$, implying that the bias is almost $200\%$. Our corrected estimator $\widehat{\beta}_{L}$ successfully reduces this bias to $0.328$, which is one third of the uncorrected and roughly $60\%$ biased. As the sample sizes increase, the bias systematically becomes smaller both for the uncorrected and corrected estimators, with dramatic improvements observed from the corrected ones both for the stiatic and dynamic models. We indeed observe that the RMSE of $\widehat{\beta}_{L}$ is twice as large as that of $\widehat{\beta}$ for the static model under $N,T=(30,30)$. However, it is the only incidence here and we attribute this to some sort of numerical coincidence [^8] Below, when we compare our method with the Jackknife under the same setting, it will be seen that the RMSE of our corrected estimator is indeed smaller than the uncorrected. See also Footnote 9 below.. It can be checked that the RMSE of our corrected estimator is no worse the uncorrected one in every other case.
Table 4: Comparisons of Uncorrected and Corrected Probit Estimates, Design 2
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.943$ $2.272$ $0.151$ $0.252$ $0.073$ $0.081$
$\widehat{\beta}_{L}$ $0.328$ $5.170$ $0.041$ $0.062$ $0.017$ $0.040$
Dynamic Model
$\widehat{\beta}$ $0.909$ $1.842$ $0.166$ $0.248$ $0.084$ $0.100$
$\widehat{\beta}_{L}^{(1)}$ $0.207$ $0.396$ $0.048$ $0.078$ $0.022$ $0.042$
$\widehat{\beta}_{L}^{(2)}$ $0.202$ $0.282$ $0.049$ $0.080$ $0.022$ $0.042$
$\widehat{\rho}$ $-0.030$ $0.162$ $-0.016$ $0.073$ $-0.011$ $0.045$
$\widehat{\rho}_{L}^{(1)}$ $-0.015$ $0.140$ $-0.005$ $0.066$ $-0.004$ $0.044$
$\widehat{\rho}_{L}^{(2)}$ $-0.009$ $0.141$ $-0.001$ $0.067$ $-0.001$ $0.045$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table 5 below reports the empirical size of the LR test for the probit model under Design 2. As before, the null hypothesis is $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model and $H_{0}:\beta_{0}=0.5$ for the static. We set the truncation parameter $\tau=0$ for the static model and $\tau=1$ for the dynamic model. Here the size distortion is much more fatal than in the logit case above. For the static model with $(N,T)=(30,30)$, the uncorrected LR test produces an empirical size of $77\%$. Even under $(N,T)=(90,90)$, the empirical size is still $72\%$, indicating a prohibitively high risk of committing a Type-I error. On the other hand, the $L(\theta)$-based LR test delivers a $14\%$ empirical size only, which is almost a $60\%$ decrease in magnitude. Similar performance is observed in other cases. The results here demonstrate that our method can reduce the type-I error risks by a major margin. Indeed, our test still over-rejects under $(N,T)=(90,90)$. This is anticipated, because the IPP under Design 2 is much more drastic and $(N,T)=(90,90)$ may not be sufficient for the test statistic to be well-approximated by $\chi^{2}$. However, it is clear that, as $(N,T)$ increase, the empirical size does become smaller. We have tried additional small-scale simulations (not reported though) with even larger $(N,T)$ and found that the empirical size still decreases with $(N,T)$. We similarly report the results from the LM and the Wald test in Appendix B.2. We report the empirical power under Design 2 in the Appendix B, since they do not add too much value in the discussion here. A quick summary for those results is that our procedure produces tests with stronger power in Design 2.
Table 5: Empirical Size of LR Tests Based on Corrected and Uncorrected Likelihoods for Logit Model, Design 2
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $77\%$ $39\%$ $71\%$ $37\%$
$(60,60)$ $76\%$ $20\%$ $66\%$ $19\%$
$(90,90)$ $72\%$ $14\%$ $63\%$ $14\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level of $5\%$. The number of replications is $1000$.
Finally, we compare our method with the Jackknife of Fernández-Val and Weidner (2016). We consider the same probit model and continue the same setting and design as above. Table 6 displays the estimation bias and the RMSE of the uncorrected estimators $\{\widehat{\rho},\widehat{\beta}\}$, our corrected estimator $\{\widehat{\rho}_{L},\widehat{\beta}_{L}\}$, and the Jackknife-corrected estimator $\{\widehat{\rho}_{J},\widehat{\beta}_{J}\}$ for the dynamic model. For the static model, only $\widehat{\beta}$, $\widehat{\beta}_{L}$, and $\widehat{\beta}_{J}$ apply and are reported. We set the truncation parameter $\tau=0$ for static model and $\tau=1$ for dynamic model. Across all pairs of $(N,T)$, the Jackknife-corrected $\widehat{\beta}_{J}$ consistently exhibits larger bias and RMSE than our $\widehat{\beta}_{L}$. Even worse, $\widehat{\beta}_{J}$ shows a larger bias and RMSE than the uncorrected $\widehat{\beta}$ across all pairs of $(N,T)$. For instance, with $(N,T)=(30,30)$ and the static model, $\widehat{\beta}_{J}$ shows a bias of $-4.929$, which is roughly $900\%$ biased (in magnitude) relative to the true value $\beta_{0}=0.5$. Observe that the uncorrected estimator $\widehat{\beta}$ only shows a roughly $170\%$ bias relative to the true value. Our corrected $\widehat{\beta}_{L}$ performs the best among the three, showing a $36\%$ bias [^9] Also notice that the RMSE of $\widehat{\beta}_{L}$ in Table 6 is much smaller than that of $\widehat{\beta}_{L}$ in Table 4, where both tables are essentially under the same setting. This motivates our attribution of the large RMSE in Table 4 to numerical coincidence.. To summarize, we find that the Jackknife is not a suitable candidate for DN-HP models. Although it does reduce the bias for the homogeneous coefficient for large $N,T$, the bias the Jackknife causes to the heterogeneous parameter is dangerously large, especially in small sample.
Table 6: Comparison with the Jackknife Bias Correction Method
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
$\widehat{\beta}_{J}$ $-4.929$ $14.383$ $-0.504$ $0.981$ $-0.105$ $0.214$
Dynamic Model
$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
$\widehat{\beta}_{J}$ $-3.352$ $7.708$ $-0.558$ $1.125$ $-0.118$ $0.217$
$\widehat{\rho}$ $-0.046$ $0.169$ $-0.022$ $0.072$ $-0.014$ $0.046$
$\widehat{\rho}_{L}$ $-0.028$ $0.145$ $-0.011$ $0.067$ $-0.009$ $0.044$
$\widehat{\rho}_{J}$ $-0.029$ $0.169$ $-0.006$ $0.068$ $-0.002$ $0.044$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The true value is $(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\beta_{0}=0$ for the static model. Accordingly, $(\widehat{\beta},\widehat{\rho})^{\prime}$ (or $\widehat{\beta}$) denotes the uncorrected estimator; $(\widehat{\beta}_{L},\widehat{\rho}_{L})^{\prime}$ (or $\widehat{\beta}_{L}$) is our bias-corrected MLE with truncation parameter $\tau=1$ for the dynamic model or $\tau=0$ for the static model; and $(\widehat{\beta}_{J},\widehat{\rho}_{J})^{\prime}$ (or $\widehat{\beta}_{J}$) represents the Jackknife bias-corrected estimators. The number of replications is $1000$.
5 Empirical Analysis
In this section we apply our likelihood-based bias correction method to examining the determinants of the labor force participation (LFP) decision of single mothers. In particular, we look at the impact of the number of children on the decision of the mother to engage in paid employment.
An extensive literature in labor economics has studied the labor supply decisions of married women (Killingsworth and Heckman, 1986; Angrist and Evans, 1998; Blau and Kahn, 2007). These studies have uncovered the impacts of a variety of economic variables, including female market wage and husband income (Mincer, 1962), education (Heath and Jayachandran, 2018), childcare costs (Connelly, 1992), the cost of home technology (Greenwood et al., 2016), and culture norms (Fernández, 2013). Among these variables, the number of children consistently emerges as one of the most important determinants of female labor supply (Nakamura and Nakamura, 1992). This is unsurprising since women continue to bear a disproportionate share of child-rearing responsibilities (Aguero and Marks, 2008).
In contrast to the substantial body of research on the labor supply behavior of married women, the labor supply decisions of single mothers have received relatively limited attention, with only a few exceptions (Kimmel, 1998; Blundell et al., 2016). Single mothers, however, face distinctive challenges when it comes to balancing work and child-rearing responsibilities due to the absence of a second earner in the household. Moreover, their employment decisions may have a more pronounced impact on their children’s well-being compared to the decisions of married women and should thus be of great importance to economists and policymakers.
To study the labor supply decisions of single mothers, we compile a data set from waves $20-30$ of the Panel Study of Income Dynamics (PSID), which span the period of 1987 to 1997. Our sample includes only single mothers, defined as unmarried female household heads with at least one child. The dependent variable is labor force participation status, defined as whether the individual worked any hours during the interview year. Our main explanatory variable is the number of children under 18 in the household. Following Dhaene and Jochmans (2015), we restrict our sample to individuals aged $18-60$ whose LFP status changed at least once during the period. The final sample comprises $N=90$ single mothers whose labor supply status is continuously observed over $T=10$ years from 1987 to 1996. Further details on data construction are provided in Appendix C. On average, 27 percent of these individuals switched in and out of the labor force in each year (Figure C.1--C.2). Table C.1 presents the summary statistics.
Let $Y_{it}\in\left\{0,1\right\}$ denote the labor force participation status of single mother $i$ in year $t$. Let $X_{it}$ denote her number of children. To assess the impact of $X_{it}$ on $Y_{it}$, we estimate the following two logit regressions:
$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\beta X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\}$ (5.1)
$Y_{it}$ $=\mathds{1}\left\{\rho Y_{it-1}+\left(\beta+\alpha_{i}+\gamma_{t}\right)X_{it}+\mu_{i}+\varphi_{t}+\varepsilon_{it}>0\right\},$ (5.2)
where $\varepsilon_{it}$ follows a standard logistic distribution, and $\sum_{i}\mu_{i}=\sum_{i}\alpha_{i}=\sum_{i}\gamma_{t}=0$. Model (5.1) is a dynamic homogeneous slope model where $\beta$ represents the average effect of an additional child on the log odds of a single mother’s LFP. Model (5.2) is a dynamic heterogeneous slope model, where $\beta_{it}=\beta+\alpha_{i}+\gamma_{t}$ captures the heterogeneous impact $X_{it}$ may have on the log odds of $Y_{it}$. Model (5.2) nests model (5.1) as a special case when $\beta_{it}$ does not depend on $i$ and $t$.
As discussed in section 1 , when the true impact of $X_{it}$ varies among the population, model (5.1) is appropriate only if $\beta_{it}$ and $X_{it}$ are uncorrelated. In practice, they may correlate due to self-selection. For instance, single mothers who choose to have more children might be those with sufficient financial resources or access to childcare services, such that their labor force participation is less affected by additional children. Alternatively, a correlation between $\beta_{it}$ and $X_{it}$ could arise from common time trends. As we track single mother households from 1987 to 1996, more children were born while factors such as rising childcare and schooling costs, improvements in home technology, and evolving welfare policies changed the impact of each additional child on their mother’s LFP. In both scenarios, model (5.2) would be the appropriate model to use as it controls for unobserved heterogeneity in $\beta_{it}$ by incorporating individual and time fixed effects into the slope and allowing them to correlate with $X_{it}$.
Given our small sample size, direct estimation of both models could suffer from severe incidental parameter problems, which we address using our bias correction procedure. Table 7 presents the estimation results. Panel A reports results for model (5.1) under the homogeneous slope assumption. For each parameter, we provide the MLE and bias-corrected estimates, along with their standard errors. Examining the bias-corrected estimates, we observe that an increase in the number of children appears to reduce a single mother’s likelihood of labor force participation, while past LFP status strongly predicts current LFP. However, this relationship shifts dramatically once we account for slope heterogeneity. Panel B presents the results for model (5.2), allowing for heterogeneous slopes. Here, the bias-corrected estimates reveal a positive average impact of additional children on a mother’s labor supply: each additional child on average increases the log odds of labor force participation by 0.126, which corresponds to an average increase of 5.5 percent in the likelihood of employment. This indicates that, on average, single mothers with more children are indeed more likely to work, likely due to the financial demands of supporting multiple children independently. Additionally, unlike the homogeneous slope model, the serial correlation in labor force participation remains statistically significant but diminishes in magnitude, suggesting that what appears to be strong state dependence in LFP under the homogeneous slope model may in part be driven by unobserved heterogeneity in how children affect labor supply decisions across individuals. Together, these findings underscore the importance of controlling for unobserved slope heterogeneity, as failing to do so can lead to substantially different conclusions.
In addition to highlighting the difference between homogeneous and heterogeneous slope models, Table 7 demonstrates the importance of bias-correction in estimating dynamic nonlinear panel data models. Examining the results in Panel B, we observe that the MLE overestimates the impact of the number of children while underestimating the degree of serial correlation in LFP. Bias correction results in an 18 percent lower estimate of the former and a 57 percent higher estimate of the latter. Moreover, the corrected estimates yield smaller standard errors, improving statistical significance across all parameters.
Finally, in Figure 2, we plot the distribution of the estimated $\beta_{it}$ from model (5.2), with summary statistics in Table C.2. Each $\beta_{it}$ estimate is constructed as the sum of the bias-corrected common effect $\beta$ and the estimated individual and time effects $\alpha_{i}$ and $\gamma_{t}$. These estimates reveal substantial heterogeneity in how the number of children influences single mothers’ labor force participation. While the mean effect is positive, over half of the individual-time observations show a decrease in LFP with each additional child, suggesting that the average effect alone may inadequately capture the diverse impacts across individuals and over time. This reinforces the importance of accounting for both individual and temporal variations in slope coefficients, as the impact of children on labor force participation can vary significantly based on unobserved factors unique to each mother and time period.
In conclusion, our analysis demonstrates the efficacy of our likelihood-based bias correction method for estimating dynamic nonlinear models of labor supply with multidimensional heterogeneities. We show the importance of controlling for unobserved slope heterogeneity when it may correlate with regressors and the value of panel data models that accommodate two-way fixed effects in both the intercept and the slope. Our findings reveal that the number of children has, on average, a positive influence on a single mother’s labor force participation, but such effect is highly heterogeneous across individuals and time periods. This heterogeneity underscores the complex nature of labor supply decisions and the unique challenges each single mother faces in balancing work and child-rearing responsibilities.
Table 7: Single Mother Labor Force Participation: Estimation Results
         Number of Children          Lagged Participation
Uncorrected Corrected Uncorrected Corrected
                Panel A. Homogeneous Slope
Model parameter -0.954 -0.880 0.584 0.535
(0.240) (0.218) (0.113) (0.103)
                Panel B. Heterogeneous Slope
Model parameter 0.153 0.126 0.051 0.080
(0.021) (0.000) (0.020) (0.000)
  • Notes: Standard errors in parentheses. Panel A reports the estimation results of the homogeneous slope model. Panel B reports the estimation results of the heterogeneous slope model. The “uncorrected” and the “corrected” columns show respectively the MLE and the bias-corrected parameter estimates. Data source: PSID 1987 - 1997.
Figure 2: Heterogeneous Impact of Children on Single Mother Labor Force Participation
6 Final Remarks
In this paper, we propose a likelihood-based analytical bias correction procedure for a general class of dynamic nonlinear heterogeneous parameter models subject to the incidental parameter problem. We give the analytical form of a corrected likelihood and show that such a corrected likelihood delivers point estimates that are asymptotically unbiased and test statistics that are asymptotically $\chi^{2}$-distributed. Simulation studies and empirical analyses support our claims.
Several issues are, however, not formally addressed in this paper. We conjecture, but have not formally investigated, that the average partial effects associated these models are asymptotically unbiased. Specifically, we conjecture that the rate of convergence of the average partial effects are only $O(1/\sqrt{T})$ or $O(1/\sqrt{N})$, slower than the potential bias, indicating that bias corrections may have no benefit in the limit. This result is formally shown by Fernández-Val and Weidner (2016) for models with additive fixed effects in the intercept. Second, we have not applied our procedure to general M-estimation models. We intuitively believe that our approach applies to them as long as the objective function is differentiable, only with minor modifications regarding the information matrix equality. We would like to dedicate these extensions to future research.
Acknowledgement
We would like to express our gratitude to valuable comments and suggestions from Geert Dhaene, Yanqin Fan, Jinyong Hahn, Chen Hsiao, Whitney Newey, and Jun Yu. Our research assistant Yuchi Liu helped us with various tasks for the empirical analysis. Leng’s research is partially supported by the National Natural Science Foundation of China Grant 72001183. Sun acknowledges the support from the National Natural Science Foundation of China under Grant Number 72203032.
Appendix
[…]
AppendixAppendix
A Proofs
In this appendix, we prove the asymptotic properties of $L(\theta)$, the corrected likelihood, $\widehat{\theta}_{L}$ the corrected estimator, and $\{\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$, the three test statistics we consider in this paper.
To continue, we first introduce three Lemmas. These lemmas are useful in establishing Theorems 1 (asymptotic bias expansion) and 2 (corrected likelihood). The proof of Corollaries 1 (estimators) and 2 (tests) then follow.
Define the partition
$H(\theta)=:\begin{pmatrix}H_{\alpha\alpha^{\prime}}\left(\theta\right),&H_{\alpha\gamma^{\prime}}\left(\theta\right)\\ H_{\gamma\alpha^{\prime}}\left(\theta\right),&H_{\gamma\gamma^{\prime}}\left(\theta\right)\end{pmatrix},$
where $H_{\alpha\alpha^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(N-1\right)K$, $H_{\gamma\gamma^{\prime}}\left(\theta\right)$ is $\left(T-1\right)K\times\left(T-1\right)K$, $H_{\alpha\gamma^{\prime}}\left(\theta\right)$ is $\left(N-1\right)K\times\left(T-1\right)K$, and $H_{\gamma\alpha^{\prime}}\left(\theta\right)=\left[H_{\alpha\gamma^{\prime}}\left(\theta\right)\right]^{\prime}$. The next lemma states that the quadratic form $\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}$ can be decomposed into two dominating terms of order $O(1/\sqrt{NT})$. They establish the two additive bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, arising from estimating individual effects and time effects, respectively.
Lemma A.1
Under Assumption 1, we have
$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E\{}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\}+o(1/\sqrt{NT}),$
where $o(\cdot)$ is uniform in $\theta$, and
$\sup_{\theta\in\Theta}\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}=O\left(1/T\right),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$
Proof. Seeing $\mathbb{E}s(\theta)=0$, the claim without $\mathbb{E}$ is
$\quad\quad s^{\prime}(\theta)[-{H}(\theta)]^{-1}s\left(\theta\right)$
$=[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]$
$=[s(\theta)-\mathbb{E}s(\theta)]^{\prime}[-\mathbb{E}{H}\left(\theta\right)]^{-1}[s(\theta)-\mathbb{E}s(\theta)]^{\prime}(1+o_{\mathbb{P}}(1))$
$=-\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)+2\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right](1+o_{\mathbb{P}}(1)).$ (A.1)
Since $K$ is fixed, we assume without loss of generality that $K=1$. In what follows, we show
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]=O\left(1/(NT)\right),$ (A.2)
and
$\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]=O(1/T),\qquad\sup_{\theta\in\Theta}\mathbb{E}\left[\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right]=O(1/N).$ (A.3)
Thus, $\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}\left(\theta\right)\right]$ is negligible in the expansion relative to the other two terms from the quadratic form.
First, under Assumptions 1 and 1, and by Davydov’s inequality, it follows that $\sup_{i,j}\sup_{\theta\in\Theta}\operatorname*{Var}[h_{i,j}^{\alpha}(\theta)]=o(1)$ as $N,T\rightarrow\infty$, where $\operatorname*{Var}$ denotes the variance and $h_{i,j}^{\alpha}(\theta)$ is the $\left(i,j\right)$-th element of $N\cdot H_{\alpha\alpha^{\prime}}\left(\theta\right)$ with $i,j=1,\ldots,N-1$. Invoking the Markov’s inequality,
$\sup_{i,j}\sup_{\theta\in\Theta}\left\{h_{i,j}^{\alpha}(\theta)-\mathbb{E}[h_{i,j}^{\alpha}(\theta)]\right\}=o_{\mathbb{P}}(1).$
Similarly, we have
$\sup_{t,s}\sup_{\theta\in\Theta}\left\{h_{t,s}^{\gamma}(\theta)-\mathbb{E}[h_{t,s}^{\gamma}(\theta)]\right\}$ $=o_{\mathbb{P}}(1),$
$\sup_{i,t}\sup_{\theta\in\Theta}\left\{h_{i,t}^{\alpha\gamma}(\theta)-\mathbb{E}[h_{i,t}^{\alpha\gamma}(\theta)]\right\}$ $=o_{\mathbb{P}}(1),$
where $h_{t,s}^{\gamma}(\theta)$ denote the $(t,s)$-th element of $T\cdot H_{\gamma\gamma^{\prime}}(\theta)$, with $t,s=1,\ldots,T-1$, and $h_{i,t}^{\alpha\gamma}(\theta)$ the $(i,t)$-th element of $NT\cdot H_{\alpha\gamma^{\prime}}(\theta)$, for $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$. Therefore, we have
$N\left\|H_{\alpha\alpha^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\alpha^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),\qquad T\left\|H_{\gamma\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\gamma\gamma^{\prime}}(\theta)]\right\|_{\max}=o_{\mathbb{P}}(1),$
$NT\left\|H_{\alpha\gamma^{\prime}}(\theta)-\mathbb{E}[H_{\alpha\gamma^{\prime}}(\theta)]\right\|_{\max}$ $=o_{\mathbb{P}}(1),$
uniformly for $\theta\in\Theta$, where $\|A\|_{\max}$ denotes the max norm of a matrix $A$. In addition, by matrix inversion in block form, it follows that
$\sup_{\theta\in\Theta}\left\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(T),\qquad\sup_{\theta\in\Theta}\left\|H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\right\|_{\max}=O(N),$
$\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}$ $=O(1).$
For $i=1,\ldots,N-1$ and $t=1,\ldots,T-1$, define $h_{i,t}^{\ast\alpha\gamma}(\theta)$ as the $(i,t)$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}(\theta)$. We first show (A.2). We have
$\quad\quad\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}\left(\theta\right))H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)(\sqrt{N}T\widetilde{s}_{\gamma}(\theta))\right\}$
$=\frac{1}{\sqrt{NT}}\mathbb{E}\left\{\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\cdot\left[\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)-\widetilde{s}_{N,s}^{\alpha}(\theta)\right]\left[\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right\}$
$=\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\mathbb{E}\left[\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right]$
$\quad\quad-\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left[\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\widetilde{s}_{j,T}^{\gamma}(\theta)\right)\right]$
$\quad\quad-\frac{1}{\sqrt{NT}}\mathbb{E}\left\{\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\right)\left[\sum_{t=1}^{T-1}\widetilde{s}_{N,t}^{\gamma}(\theta)\left(\sum_{i=1}^{N-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right]\right\}$
$\quad\quad+\frac{1}{\sqrt{NT}}\left(\sum_{i=1}^{N-1}\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\widetilde{s}_{N,T}^{\gamma}(\theta)\right)$
$=:A_{NT,1}(\theta)-A_{NT,2}(\theta)-A_{NT,3}(\theta)+A_{NT,4}(\theta).$ (A.4)
By Assumptions 1 and 1, and using Davydov’s inequality, we obtain for some $\eta>2$ that
$\left|A_{NT,1}(\theta)\right|$ $=\frac{1}{\sqrt{NT}}\sum_{i=1}^{N-1}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T-1}\widetilde{s}_{i,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right)+\sum_{t=1}^{T-1}\widetilde{s}_{i,T}^{\alpha}(\theta)\widetilde{s}_{i,t}^{\gamma}(\theta)h_{i,t}^{\ast\alpha\gamma}(\theta)\right]\right|$
$\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
×$\quad\quad\times\sup\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT}).$
Similarly,
$\left|A_{NT,2}(\theta)\right|$ $=\frac{T}{\sqrt{NT}}\sum_{i=1}^{N-1}\left(\frac{1}{T}\sum_{t=1}^{T-1}\left|h_{i,t}^{\ast\alpha\gamma}(\theta)\right|\right)\left|\mathbb{E}\left(\sum_{s=1}^{T}\widetilde{s}_{i,s}^{\alpha}(\theta)\widetilde{s}_{i,T}^{\gamma}(\theta)\right)\right|$
$\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\left(\sup_{i}\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT});$
$\left|A_{NT,3}(\theta)\right|$ $=\frac{N}{\sqrt{NT}}\left|\mathbb{E}\left[\left(\sum_{s=1}^{T}\widetilde{s}_{N,s}^{\alpha}(\theta)\right)\left(\sum_{t=1}^{T-1}\widetilde{s}_{N,t}^{\gamma}(\theta)\left(\frac{1}{N}\sum_{i=1}^{N-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\right)\right]\right|$
$=\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)$
×$\quad\quad\times\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-2}\left[\left(1-\frac{m}{T-1}\right)[a_{i}(m)]^{1-2/\eta}\right]+\frac{1}{T-1}\sum_{m=1}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT});$
$\left|A_{NT,4}(\theta)\right|$ $\leq\sqrt{NT}\left(\sup_{\theta\in\Theta}\left\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\right\|_{\max}\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left(\sum_{m=0}^{T-1}[a_{i}(m)]^{1-2/\eta}\right)$
$=O(\sqrt{NT}).$
Therefore, (A.2) follows from (A.4).
Let $h_{i,j}^{\ast\alpha}(\theta)$ denote the $(i,j)$-th element of $T^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)$, $i,j=1,\ldots,N-1$. By Assumption 1, we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ uniformly in $\theta$. Recall, $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$ and $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$. Note that
$N\cdot H_{\alpha\alpha^{\prime}}(\theta)=D_{1}[N\cdot\partial_{\alpha\alpha^{\prime}}l(\theta,\phi)|_{\phi=\phi(\theta)}]D_{1}^{\prime},$
implying $N\cdot H_{\alpha\alpha^{\prime}}(\theta)$ can be written as the sum of a diagonal matrix and a rank-one matrix. Thus, following from Miller (1981), we have $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Now,
$\quad\quad T\mathbb{E}\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right]$
$=\frac{1}{N}\mathbb{E}\left\{(N\sqrt{T}\widetilde{s}_{\alpha}^{\prime}(\theta))[N^{-1}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)](N\sqrt{T}\widetilde{s}_{\alpha}(\theta))\right\}$
$=\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\alpha}(\theta)\mathbb{E}\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)^{2}$
$\quad\quad+\frac{2}{N}\sum_{1\leq i<j\leq N-1}h_{i,j}^{\ast\alpha}(\theta)\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right]$
$=:I_{NT,1}(\theta)+I_{NT,2}(\theta).$ (A.5)
Here, by Assumptions 1 and 1, applying Davydov’s inequality gives, for some $\eta>2$,
$\quad\quad\sup_{\theta\in\Theta}\left|I_{NT,1}(\theta)\right|$
$=\sup_{\theta\in\Theta}\frac{1}{NT}\sum_{i=1}^{N-1}\left|h_{i,i}^{\ast\alpha}(\theta)\right|\left|\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\widetilde{s}_{i,s}^{\alpha}(\theta)\right]+\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\left[\widetilde{s}_{N,t}^{\alpha}(\theta)\widetilde{s}_{N,s}^{\alpha}(\theta)\right]\right|$
$\leq\left(\frac{1}{N}\sum_{i=1}^{N-1}|h_{i,i}^{\ast\breve{\alpha}}(\theta)|\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}\left(1-\frac{m}{T}\right)[a_{i}(m)]^{1-2/\eta}\right)$
$=O\left(1\right)$
as $N,T\rightarrow\infty$, where the conditional cross-sectional independence leads to the first equality. Similarly
$\quad\quad\sup_{\theta\in\Theta}\left|I_{NT,2}(\theta)\right|$
$=\frac{2}{N}\sum_{1\leq i<j\leq N-1}\left\{\left|h_{i,j}^{\ast\alpha}(\theta)\right|\mathbb{E}\left[\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{N,t}^{\alpha}(\theta)\right)^{2}\right]\right\}$
$\leq\left(\frac{2}{N}\sum_{1\leq i<j\leq N-1}|h_{i,j}^{\ast\alpha}(\theta)|\right)\left(\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\right)\sup_{i}\left([a_{i}(0)]^{1-2/\eta}+2\sum_{m=1}^{T-1}\left(1-\frac{m}{T}\right)[a_{i}(m)]^{1-2/\eta}\right)$
$=O\left(1\right),$
noting that $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$. Thus, the first part of (A.3) follows from (A.5). The second part of (A.3) follows from a similar procedure.
Combining (A.1), (A.2), and (A.3),
$\mathbb{E}\left\{s^{\prime}(\theta)[-{H}(\theta)]^{-1}s(\theta)\right\}$ $=\mathbb{E}\left\{\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\right\}+\mathbb{E}\left\{\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right\}$
$\quad\quad+\mathbb{E}\left\{\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}\left(\theta\right)H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)\right]\cdot o_{\mathbb{P}}(1)\right\}.$
By Assumption 1, for $S\in\{0,1,2\}$, we have
$\sup_{i,t}\mathbb{E}\left(|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}\left(\theta,\phi\right)|\mathds{1}\{|\partial_{\phi_{1}\ldots\phi_{S}}l_{i,t}\left(\theta,\phi\right)|>a\}\right)\rightarrow 0$
as $a\rightarrow\infty$, which shows that $\{\partial_{\phi_{1}\ldots\phi_{S}}l_{it}\left(\theta,\phi\right)\}_{i,t}$ are uniformly integrable. By the $L_{1}$ convergence theorem, we hence have
$\sqrt{NT}\mathbb{E}\left\{\left[\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+\widetilde{s}_{\gamma}^{\prime}\left(\theta\right)H_{\gamma\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)\right]\cdot o_{\mathbb{P}}(1)\right\}\rightarrow 0,$
as $N,T\rightarrow\infty$. This completes the proof.
The next lemma shows that the remainder term in Equation (2.13), $r^{s}\left(\theta\right):=r^{s}\left(\theta,a\right)$ is $O_{\mathbb{P}}\left(1/\left(NT\right)\right)$ uniformly, where $r^{s}\left(\theta,a\right)$ is defined as [^10] Feng et al. (2014) point out that the expansion in (2.13) needs to be done component-by-component. This leads to a different $a_{j}$ for each component of $r^{s}\left(\theta,a\right)$.
$r^{s}\left(\theta,a\right)$ $:=\left(r_{\psi_{1}}^{s}\left(\theta,a_{1}\right),\ldots,r_{\psi_{(N+T-2)K}}^{s}(\theta,a_{(N+T-2)K})\right)^{\prime},$ (A.6)
$r_{\psi_{j}}^{s}\left(\theta,a_{j}\right)$ $:=\frac{1}{2}\left(\widehat{\psi}\left(\theta\right)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi(\theta)\right),$
$G_{\psi_{j}}(\theta,a_{j})$ $:=\partial_{\psi_{j}}H(\theta,a_{j}\widehat{\psi}(\theta)+(1-a_{j})\psi\left(\theta\right))$
with $a:=(a_{1},\ldots,a_{(N+T-2)K})^{\prime}$ for each $a_{j}\in(0,1)$, and $\psi_{j}$ denotes the $j$-th component of $\psi$. Here $G_{\psi_{j}}(\theta,a_{j})$ is the derivative of each element of $H(\theta,\psi)$ with respect to $\psi_{j}$, evaluated at the point $a_{j}\widehat{\psi}\left(\theta\right)+\left(1-a_{j}\right)\psi\left(\theta\right)$.
Lemma A.2
Under Assumption 1,
$\sup_{\theta\in\Theta}\sup_{a}\left\|r^{s}(\theta,a)\right\|_{\max}=O_{\mathbb{P}}\left(1/\left(NT\right)\right),$
as $N$ and $T\rightarrow\infty$, where $\sup_{a}$ is taken over $a_{j}\in(0,1)$ for all $j=1,\ldots,(N+T-2)K$.
Proof. In the proof, we retain the notation and setup from Lemma A.1. From (2.13), it is evident that the dominant term in the expansion of $\widehat{\psi}(\theta)-\psi\left(\theta\right)$ is $[-{H}(\theta)]^{-1}s(\theta)$. Specifically,
$[-{H}(\theta)]^{-1}s(\theta)$ $=[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)](1+o_{\mathbb{P}}(1))$
$=-\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}+[-\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\cdot o_{\mathbb{P}}(1)$ (A.7)
where $o_{\mathbb{P}}(1)$ is uniform in both $i$ and $t$. Let $u_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)$. It follows that
$N\sqrt{T}u_{i}^{\alpha}(\theta)=N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\left[\widetilde{s}_{j,t}^{\alpha}(\theta)-\widetilde{s}_{N,t}^{\alpha}(\theta)\right]\right)\right].$ (A.8)
Using the result from Miller (1981), we know that $\sup_{i\geq 1}h_{i,i}^{\ast\alpha}(\theta)=O(1)$ and $\sup_{i\neq j}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in $\theta$, implying $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. Since $1/\sqrt{T}\sum_{t=1}^{T}\widetilde{s}_{i,t}^{\alpha}(\theta)=O_{\mathbb{P}}(1)$ uniformly over $i$ by Assumptions 1 and 1, we obtain
$N\sum_{j=1}^{N-1}\left[h_{i,j}^{\ast\alpha}(\theta)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{j,t}^{\alpha}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad N\left(\sum_{j=1}^{N-1}h_{i,j}^{\ast\breve{\alpha}}(\theta)\right)\left(\frac{1}{\sqrt{T}}\sum_{t=1}^{T}\widetilde{s}_{N,t}^{\alpha}(\theta)\right)=O_{\mathbb{P}}(1).$
Therefore, $N\sqrt{T}u_{i}^{\alpha}(\theta)=O_{\mathbb{P}}(1)$ uniformly over $i$. Let $u_{i}^{\alpha\gamma}(\theta)$ denote the $i$-th element of $H_{\alpha\gamma^{\prime}}^{\ast}\left(\theta\right)\widetilde{s}_{\gamma}(\theta)$. Given that $\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\|_{\max}=O(1)$, we have
$T\sqrt{N}u_{i}^{\alpha\gamma}(\theta)=\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\left[\widetilde{s}_{j,t}^{\gamma}(\theta)-\widetilde{s}_{j,T}^{\gamma}(\theta)\right]\right)\right].$ (A.9)
Since $\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)=O(1)$, which follows from $\mathbb{E}[\partial_{\breve{\alpha}\breve{\gamma}^{\prime}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=\partial_{\breve{\gamma}^{\prime}}\mathbb{E}[\partial_{\breve{\alpha}}\ell(\theta,\psi)|_{\psi=\psi(\theta)}]=0$, for $\breve{\alpha}:=\left(\alpha_{1}^{\prime},\ldots,\alpha_{N-1}^{\prime}\right)^{\prime}$ and $\breve{\gamma}:=\left(\gamma_{1}^{\prime},\ldots,\gamma_{T-1}^{\prime}\right)^{\prime}$, it holds that
$\sum_{t=1}^{T-1}\left[h_{i,t}^{\ast\alpha\gamma}(\theta)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,t}^{\gamma}(\theta)\right)\right]=O_{\mathbb{P}}(1),\qquad\left(\sum_{t=1}^{T-1}h_{i,t}^{\ast\alpha\gamma}(\theta)\right)\left(\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\widetilde{s}_{j,T}^{\gamma}(\theta)\right)=O_{\mathbb{P}}(1)$
uniformly in $\theta$. From (A.8) and (A.9), we conclude that
$\sup_{\theta\in\Theta}\|H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right),\qquad\sup_{\theta\in\Theta}\|H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right).$ (A.10)
Following a similar proof, we also have
$\sup_{\theta\in\Theta}\|H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(T\sqrt{N})\right),\qquad\sup_{\theta\in\Theta}\|H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\breve{\alpha}}(\theta)\|_{\max}=O_{\mathbb{P}}\left(1/(N\sqrt{T})\right).$ (A.11)
Thus, combining (A.7), (A.10), and (A.11), we have
$\left\|[-{H}(\theta)]^{-1}s(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right),$
where $o_{\mathbb{P}}(\cdot)$ is uniform in $\theta$. Noting that
$\widehat{\psi}(\theta)-\psi(\theta)=[-{H}\left(\theta\right)]^{-1}s(\theta)(1+o_{\mathbb{P}}(1)),$
we further have
$\left\|\widehat{\psi}(\theta)-\psi(\theta)+[\mathbb{E}{H}(\theta)]^{-1}[{s}(\theta)-\mathbb{E}{s}(\theta)]\right\|_{\max}=o_{\mathbb{P}}\left(1/(NT)^{3/4}\right).$ (A.12)
From (A.12), we conclude that $\sup_{\theta\in\Theta}\|G_{\psi_{j}}\left(\theta,a_{j}\right)-\partial_{\psi_{j}}H(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Without loss of generality, let $\psi_{j}=\alpha_{i}$ for $i=1,\ldots,N-1$. We can express $\partial_{\alpha_{i}}H(\theta)$ as
$\partial_{\alpha_{i}}H(\theta)=\begin{pmatrix}\partial_{\alpha_{i}}{H}_{\alpha\alpha^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\alpha\gamma^{\prime}}(\theta)\\ \partial_{\alpha_{i}}{H}_{\gamma\alpha^{\prime}}\left(\theta\right),&\partial_{\alpha_{i}}{H}_{\gamma\gamma^{\prime}}(\theta)\end{pmatrix}.$
We then have
$N\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}\left(\theta\right)$ $=\frac{1}{T}\sum_{t=1}^{T}\operatorname*{diag}\left(0,\ldots,0,\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)},0,\ldots,0\right)$
$\quad\quad-\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\iota_{N-1}\iota_{N-1}^{\prime},$
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}\left(\theta\right)$ $=\operatorname*{diag}\left\{\partial_{\gamma_{1}\gamma_{1}\alpha_{i}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\gamma_{T-1}\gamma_{T-1}\alpha_{i}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right\}$
$\quad\quad+\partial_{\gamma_{T}\gamma_{T}\alpha_{i}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\,\iota_{N-1}\iota_{N-1}^{\prime},$
$NT\cdot\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}\left(\theta\right)$ $=\left(\partial_{\alpha_{i}\alpha_{i}\gamma_{1}}l_{i,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{i}\alpha_{i}\gamma_{T-1}}l_{i,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes e_{i}$
$\quad\quad+\left(\partial_{\alpha_{N}\alpha_{N}\gamma_{1}}l_{N,1}(\theta,\phi)|_{\phi=\phi(\theta)},\ldots,\partial_{\alpha_{N}\alpha_{N}\gamma_{T-1}}l_{N,T-1}(\theta,\phi)|_{\phi=\phi(\theta)}\right)\otimes\iota_{N-1},$
where $e_{i}$ is an $(N-1)\times 1$ vector with the $i$-th component equal to $1$ and all other elements equal to 00. Using (A.12), for each $j=1,\ldots,N-1$, we write
$\quad\quad\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=\sup_{a_{j}\in(0,1)}\frac{1}{2}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}G_{\psi_{j}}(\theta,a_{j})\left(\widehat{\psi}(\theta)-\psi\left(\theta\right)\right)$
$=\frac{1}{2}\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}^{\prime}\partial_{\alpha_{i}}H(\theta)\begin{pmatrix}H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\\ H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\end{pmatrix}\left[1+o_{\mathbb{P}}(1)\right]$
$=\frac{1}{2}\left\{\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\alpha\alpha^{\prime}}(\theta)\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right.$
$\quad\quad+\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\gamma\gamma^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)$
$\quad\quad\left.+2\left(H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)^{\prime}\partial_{\alpha_{i}}{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}(\theta)\left(H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)\right)\right\}\left[1+o_{\mathbb{P}}(1)\right]$
$=:\frac{1}{2}\left\{C_{NT,1}(\theta)+C_{NT,2}(\theta)+C_{NT,3}(\theta)\right\}\left[1+o_{\mathbb{P}}(1)\right].$ (A.13)
Let $v_{i}^{\alpha}(\theta)$ denote the $i$-th element of $H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\alpha\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$, for $i=1,\ldots,N-1$. Using (A.8) and (A.9), it can be shown that
$\quad\quad C_{NT,1}(\theta)$
$=\frac{1}{N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}\alpha_{i}\alpha_{i}}l_{i,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)-\frac{1}{N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
$=O_{\mathbb{P}}\left(1/(NT)\right),$
uniformly in $\theta$. By similar arguments, we also have
$\sup_{\theta\in\Theta}C_{NT,2}(\theta)=O_{\mathbb{P}}\left(1/(NT)^{3/2}\right),\qquad\sup_{\theta\in\Theta}C_{NT,3}(\theta)=O_{\mathbb{P}}\left(1/(NT)^{3/2}\right).$
Therefore, when $\psi_{j}=\alpha_{i}$, we have
$\quad\quad\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=-\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
$=O_{\mathbb{P}}\left(1/(NT)\right).$ (A.14)
Similarly, when $\psi_{j}=\gamma_{t}$ ($t=1,\ldots,T-1$),
$\quad\quad\sup_{\theta\in\Theta}\sup_{a_{j}\in(0,1)}r_{\psi_{j}}^{s}(\theta,a_{j})$
$=-\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)+o_{\mathbb{P}}\left(1/(NT)\right)$
$=O_{\mathbb{P}}\left(1/(NT)\right),$ (A.15)
where $v_{t}^{\gamma}(\theta)$ denotes the $t$-th element of $H_{\gamma\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)+H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)$ for $t=1,\ldots,T-1$. This leads to the desired result.
The next lemma states that the final remainder, $r(\theta)$ in Equation (2.14), is $O\left(1/\left(NT\right)\right)$ uniformly in $\theta$, where
$r(\theta)$ $:=\frac{1}{2}(r^{s}(\theta,a))^{\prime}[H\left(\theta\right)]^{-1}r^{s}(\theta,a)+r^{l}(\theta,a_{0}),$ (A.16)
$r^{l}\left(\theta,a_{0}\right)$ $:=\frac{1}{3}\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\times\iota_{(N+T-2)K})$ (A.17)
for some $a_{0}\in(0,1)$. Here $r^{l}\left(\theta,a_{0}\right)=:r^{l}(\theta)$ as in Equation (2.12) from the main text. The dependency of $r(\theta)$ on $a$ and $a_{0}$ is suppress for simplicity here.
Lemma A.3
Under Assumption 1, $r(\theta)$ defined by (A.16) satisfies
$\sup_{\theta\in\Theta}\mathbb{E}[r(\theta)]=O\left(1/\left(NT\right)\right),$
as $N$ and $T\rightarrow\infty$.
Proof. In the proof, we retain the notation and setup from Lemmas A.1 and A.2. From Lemma A.2, we know that $NT\cdot r^{s}\left(\theta,a\right)$ is comparable to $(NT)^{3/4}\cdot s(\theta)$. Following similar arguments as in Lemma A.1, we can show that
$\mathbb{E}\left\{(r^{s}(\theta,a))^{\prime}[{H}\left(\theta\right)]^{-1}r^{s}(\theta,a)\right\}=O\left(1/(NT)\right).$
Moreover, from (A.12), (A.14), and (A.14), we have
$\quad\quad\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}(\theta,a_{0}\,\iota_{N+T-2})$
$=\frac{1}{2N}\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)\right)^{2}\left(\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{N}\alpha_{N}\alpha_{N}}l_{N,t}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
$\quad\quad+\frac{1}{2T}\left(\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)^{2}\left(\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{T}\gamma_{T}\gamma_{T}}l_{i,T}(\theta,\phi)|_{\phi=\phi(\theta)}\right)$
$\quad\quad+\left(\sum_{i=1}^{N-1}v_{i}^{\alpha}(\theta)+\sum_{t=1}^{T-1}v_{t}^{\gamma}(\theta)\right)\cdot o_{\mathbb{P}}\left(1/(NT)\right)$
$=O_{\mathbb{P}}\left(1/(NT)^{5/4}\right).$ (A.18)
By Assumption 1 and applying the $L_{1}$ convergence theorem to A.18, we find that
$\sup_{\theta\in\Theta}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=o\left(1/(NT)\right).$
Thus,
$\mathbb{E}\left[r(\theta)\right]=\frac{1}{2}\mathbb{E}\left[(r^{s}\left(\theta,a\right))^{\prime}[{H}(\theta)]^{-1}r^{s}\left(\theta,a\right)\right]+\frac{1}{3}\mathbb{E}\left[\left(\widehat{\psi}(\theta)-\psi(\theta)\right)^{\prime}r^{s}\left(\theta,a_{0}\,\iota_{N+T-2}\right)\right]=O\left(1/(NT)\right)$
uniformly in $\theta$.
Next, we prove Theorem 1 and 2 below.
Proof of Theorem 1. Following the expansion in (2.12) and (2.13), we have
$\mathbb{E[}\widehat{l}(\theta)-l(\theta)]=-\frac{1}{2}\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}+\mathbb{E}r(\theta)$
for $r\left(\theta\right)$ defined in (A.16). The claim in Theorem 1 follows directly from Lemmas A.1 and A.3.
The proof of Theorem 2 is as follows.
Proof of Theorem 2. We retain the notation introduced in Lemmas A.1-A.3. We first show
$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.19)
To show the first part of (A.19), note that
$\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)=\frac{1}{2}\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}+\frac{1}{2}\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\},$
where
$S_{\alpha\alpha}\left(\theta\right)$ $=S_{\alpha\alpha}^{1}\left(\theta\right)+S_{\alpha\alpha}^{2}\left(\theta\right),$
$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\mathbb{E}\left[s_{1}^{\alpha}\left(\theta\right)\left(s_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\right],\ldots,\mathbb{E}\left[s_{N-1}^{\alpha}\left(\theta\right)\left(s_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right]\right\},$
$S_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\left[s_{N}^{\alpha}\left(\theta\right)\left(s_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right];$
$\widehat{S}_{\alpha\alpha}\left(\theta\right)$ $=\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)+\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right),$
$S_{\alpha\alpha}^{1}\left(\theta\right)$ $:=\operatorname*{diag}\left\{\widehat{s}_{1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime},\ldots,\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\right\},$
(${}_{\alpha\alpha}^{2}\left(\theta\right)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\left[\widehat{s}_{N}^{\alpha}\left(\theta\right)\left(\widehat{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\right].$
Here we have
$\quad\quad N^{2}\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]$
$=\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}$
×$\quad\quad\times\operatorname*{diag}\left\{\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{1,t}^{\alpha}(\theta)\left(\widehat{s}_{1,s}^{\alpha}(\theta)\right)^{\prime},\ldots,\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}-\widehat{s}_{N-1,t}^{\alpha}(\theta)\left(\widehat{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right\}$
$\quad\quad+\frac{1}{T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}$
×$\quad\quad\times\operatorname*{diag}\left\{\mathbb{E}\left[\widetilde{s}_{1,t}^{\alpha}(\theta)\left(\widetilde{s}_{1,s}^{\alpha}(\theta)\right)^{\prime}\right],\ldots,\mathbb{E}\left[\widetilde{s}_{N-1,t}^{\alpha}(\theta)\left(\widetilde{s}_{N-1,s}^{\alpha}(\theta)\right)^{\prime}\right]\right\}.$ (A.20)
Under Assumptions 1 and 1, by Davydov’s inequality,
$\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\left(\widetilde{s}_{i,s}^{\alpha}(\theta)\right)^{\prime}\right]\leq 8[a_{i}(|t-s|)]^{1-2/\eta}\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta},$
where we assume without loss of generality that $K=1$. Thus, as $T\rightarrow\infty$,
$\quad\quad\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}\mathbb{E}\left[\widetilde{s}_{i,t}^{\alpha}(\theta)\left(\widetilde{s}_{i,s}^{\alpha}(\theta)\right)^{\prime}\right]$
$\leq 8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\left(\frac{1}{T}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|>\tau\right\}[a_{i}(|t-s|)]^{1-2/\eta}\right)$
$\leq 8\sup_{i,t}\{\mathbb{E}[g^{\eta}(w_{it})]\}^{2/\eta}\left(\frac{T-\tau}{T}\sum_{m=\tau+1}^{\infty}[a_{i}(m)]^{1-2/\eta}\right)$
$=o(1)$
given that $\tau/T\rightarrow 0$. Combining this with (A.20), we have
$\sup_{\theta\in\Theta}N^{2}\left\|S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right\|_{\max}=o_{\mathbb{P}}(1/T).$ (A.21)
From the proof of Lemma A.2, we also have $\sup_{\theta\in\Theta}\|\widehat{\psi}(\theta)-\psi(\theta)\|_{\max}=o_{\mathbb{P}}(1)$. Moreover, under the conditional strong mixing assumption, each element of $T^{-1}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ converges in probability to the corresponding element in $T^{-1}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ by the law of large numbers. Therefore (A.21) further leads to
$\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
$=\frac{T}{N^{2}}\operatorname*{trace}\left\{T^{-1}\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\cdot N^{2}\left[S_{\alpha\alpha}^{1}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{1}\left(\theta\right)\right]\right\}$
$=\frac{T}{N}\frac{1}{N}\sum_{i=1}^{N-1}h_{i,i}^{\ast\alpha}(\theta)\cdot o_{\mathbb{P}}(1/T)$
$=o_{\mathbb{P}}(1/T).$
A similar procedure can be followed to show
$\quad\quad\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}\left(\theta\right)\left[S_{\alpha\alpha}^{2}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)\right]\right\}$
$=\frac{T}{N^{2}}\operatorname*{trace}\left\{T^{-1}\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\cdot N^{2}\left[S_{\alpha\alpha}^{2}\left(\theta\right)-\widehat{S}_{\alpha\alpha}^{2}\left(\theta\right)\right]\right\}$
$=\frac{T}{N^{2}}\sum_{i=1}^{N-1}\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)\cdot o_{\mathbb{P}}(1/T)$
$=o_{\mathbb{P}}(1/N^{2}),$
where the last equality follows from, as shown in Lemma A.2, $\sum_{j=1}^{N-1}h_{i,j}^{\ast\alpha}(\theta)=O(1/N)$ uniformly in both $i$ and $\theta$. The above leads to
$\operatorname*{trace}\left\{\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)\left[S_{\alpha\alpha}\left(\theta\right)-\widehat{S}_{\alpha\alpha}\left(\theta\right)\right]\right\}=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.22)
Second, from Theorem 1, we also have
$\operatorname*{trace}\left\{\left[\widehat{H}_{\alpha\alpha^{\prime}}^{\ast}(\theta)-H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\right]S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot\operatorname*{trace}\left\{H_{\alpha\alpha^{\prime}}^{\ast}(\theta)S_{\alpha\alpha}\left(\theta\right)\right\}=o_{\mathbb{P}}(1)\cdot B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$ (A.23)
Combining A.22 and A.23 leads to the first part of A.19. The second part of A.19 can be shown similarly.
Noting that
$l(\theta)=\widehat{l}(\theta)+\frac{1}{2}s_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)s_{\alpha}(\theta)+\frac{1}{2}s_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)s_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$
as implied by the proof of Theorem 1, we have
$L(\theta)-l(\theta)$ $=\left[\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)\right]+\left[\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)\right]$
$\quad\quad-\left[\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)\right]-\left[\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)\right]$
$\quad\quad+o_{\mathbb{P}}(1/\sqrt{NT}).$
As from Lemma A.1, we have
$\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$
The conclusion follows by combining this and (A.19).
We continue to establishing the asymptotitc properties of $\{\widehat{\theta}_{L},\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$. We first prove Corollary 1 below which states the consistency and asymptotic normality of $\widehat{\theta}_{L}$.
Proof of Corollary 1. Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}$:
$0=\triangledown_{\theta}L(\widehat{\theta}_{L})=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\theta_{0}$. From Theorem 2, we further have
$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (A.24)
Under Assumption 2, $\theta_{0}$ is the unique maximizer of $\mathbb{E}[l(\theta)]$. Since $L(\theta)-\mathbb{E}[l(\theta)]=o_{\mathbb{P}}(1)$, under Assumption 2, we conclude that
$\widehat{\theta}_{L}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}.$
This also implies $\overline{\theta}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}$. By Assumption 2, it follows that
$\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]=o_{\mathbb{P}}(1).$
Thus, using Assumption 2 and Slutsky’s theorem, we obtain from (A.24) that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.25)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.25) leads to
$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (A.26)
as $N,T\rightarrow\infty$, where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0$ and
$\Sigma:=-\mathbb{E}\left[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})\right],\qquad\Omega:=\left(NT\right)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (A.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).
To show the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$, we find the following lemma useful.
Lemma A.4
Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\rightarrow\infty,$
$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
where
$\Xi(\theta_{0}):=\left\{\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}[J(\theta_{0})]^{\prime}\left\{J(\theta_{0})\left(\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right)^{-1}[J(\theta_{0})]^{\prime}\right\}^{-1}J(\theta_{0})\left\{\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}.$
Proof. Consider the constrained maximization problem:
$\left\{\widehat{\theta}_{R},\widehat{\lambda}\right\}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\left[L(\theta)+\lambda^{\prime}R(\theta)\right]$
where $\lambda$ is the Lagrangian multiplier. The first-order conditions for this constrained maximization are
$\sqrt{NT}\triangledown_{\theta}L(\widehat{\theta}_{R})+[J(\widehat{\theta}_{R})]^{\prime}\sqrt{NT}\widehat{\lambda}$ $=0,$ (A.27)
$\sqrt{NT}R(\widehat{\theta}_{R})$ $=0.$ (A.28)
Following a similar proof as in Corollary 1, we have $\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=O_{\mathbb{P}}\left(1\right)$. Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ is
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{R}-\theta_{0})$
$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})](\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
where each component of $\overline{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\theta_{0}.$ Combining this with (A.27) yields that
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.29)
Under $H_{0}$ and from (A.28), we also obtain
00 0 $=\sqrt{NT}R(\widehat{\theta}_{R})$
$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1)$
$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (A.30)
Now, combining (A.29) and (A.30), we have
$\begin{pmatrix}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]&\left[J(\theta_{0})\right]^{\prime}\\ J(\theta_{0})&0\end{pmatrix}\begin{pmatrix}\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})\\ \sqrt{NT}\widehat{\lambda}\end{pmatrix}=\begin{pmatrix}-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\\ 0\end{pmatrix}+o_{\mathbb{P}}(1)$
which has the solution
$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.31)
The conclusion follows by subtracting (A.25) in Corollary 1 from both sides of (A.31).
Finally, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
Proof of Corollary 2. Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma A.4. We first show that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{LM}=\widehat{\xi}+o_{\mathbb{P}}(1),\qquad\widehat{\xi}_{Wald}=\widehat{\xi}+o_{\mathbb{P}}(1).$ (A.32)
A mean value expansion of $L(\widehat{\theta}_{R})$ around $\widehat{\theta}_{L}$ gives
${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})(\widehat{\theta}_{R}-\widehat{\theta}_{L})+\frac{1}{2}(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L})$
where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Noting that $\triangledown_{\theta}{L}(\widehat{\theta}_{L})=0$, we obtain
$\widehat{\xi}_{LR}=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L}).$
Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma A.4, we further conclude that
$\widehat{\xi}_{LR}=\widehat{\xi}+o_{\mathbb{P}}(1).$
For $\widehat{\xi}_{LM},$ note that
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\ddot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L})$
$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}(1/\sqrt{NT}),$
[…]
$\widehat{\xi}_{LM}$ $=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
$=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
$=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
For $\widehat{\xi}_{Wald}$, a mean value expansion of $R(\widehat{\theta}_{L})$ gives
$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
where each component of $\dddot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\widehat{\theta}_{R}$. Since both $\widehat{\theta}_{L}$ and $\widehat{\theta}_{R}$(under $H_{0}$) are consistent estimators for $\theta_{0},$ we obtain that
$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\dddot{\theta})^{\prime}\left\{J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right\}^{-1}J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
$=\widehat{\xi}+o_{\mathbb{P}}\left(1\right).$
Next, since $\Sigma^{-1/2}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K})$, applying the delta method delivers $\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right)$. The desired results follows from here.
B Extra Simulation Results
B.1 Logit Model
Table B.1: Empirical Size of LM and Wald Tests for Logit Model, Design 1
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $15\%$ $9\%$ $18\%$ $10\%$
$(60,60)$ $15\%$ $5\%$ $17\%$ $7\%$
$(90,90)$ $18\%$ $6\%$ $16\%$ $5\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$, Design 1
(a) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $48\%$ $32\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $40\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
(b) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $44\%$ $32\%$ $22\%$ $10\%$ $14\%$ $21\%$
$(60,60)$ $86\%$ $62\%$ $37\%$ $25\%$ $50\%$ $74\%$
$(90,90)$ $100\%$ $93\%$ $65\%$ $50\%$ $84\%$ $98\%$
Dynamic Model
$(30,30)$ $47\%$ $31\%$ $20\%$ $18\%$ $28\%$ $43\%$
$(60,60)$ $94\%$ $73\%$ $39\%$ $45\%$ $75\%$ $94\%$
$(90,90)$ $100\%$ $98\%$ $76\%$ $78\%$ $98\%$ $100\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Table B.3: Comparisons of Uncorrected and Corrected Logit Estimates, Design 2
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.454$ $1.201$ $0.100$ $0.132$ $0.060$ $0.071$
$\widehat{\beta}_{L}$ $0.101$ $0.199$ $0.020$ $0.056$ $0.010$ $0.035$
Dynamic Model
$\widehat{\beta}$ $0.527$ $1.583$ $0.109$ $0.212$ $0.059$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.117$ $0.204$ $0.024$ $0.058$ $0.009$ $0.038$
$\widehat{\beta}_{L}^{(2)}$ $0.126$ $0.278$ $0.025$ $0.059$ $0.009$ $0.041$
$\widehat{\rho}$ $-0.110$ $0.241$ $-0.055$ $0.110$ $-0.037$ $0.071$
$\widehat{\rho}_{L}^{(1)}$ $-0.041$ $0.192$ $-0.017$ $0.090$ $-0.010$ $0.059$
$\widehat{\rho}_{L}^{(2)}$ $-0.035$ $0.196$ $-0.012$ $0.091$ $-0.007$ $0.059$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table B.4: Empirical Size of LR, LM, and Wald Tests for Logit Model, Design 2
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $47\%$ $20\%$ $43\%$ $18\%$
$(60,60)$ $43\%$ $10\%$ $38\%$ $9\%$
$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $42\%$ $18\%$ $38\%$ $16\%$
$(60,60)$ $42\%$ $9\%$ $37\%$ $9\%$
$(90,90)$ $41\%$ $9\%$ $34\%$ $7\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $32\%$ $15\%$ $27\%$ $15\%$
$(60,60)$ $41\%$ $9\%$ $36\%$ $8\%$
$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.5: Empirical Power of LR, LM, and Wald Tests for Logit Model using $L(\theta)$, Design 2
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $79\%$ $62\%$ $45\%$ $12\%$ $16\%$ $26\%$
$(60,60)$ $100\%$ $95\%$ $73\%$ $40\%$ $75\%$ $93\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $75\%$ $58\%$ $41\%$ $19\%$ $26\%$ $35\%$
$(60,60)$ $100\%$ $94\%$ $69\%$ $44\%$ $80\%$ $96\%$
$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $77\%$ $60\%$ $42\%$ $12\%$ $17\%$ $28\%$
$(60,60)$ $99\%$ $95\%$ $72\%$ $41\%$ $77\%$ $93\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $82\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $73\%$ $56\%$ $38\%$ $19\%$ $27\%$ $36\%$
$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $80\%$ $94\%$
$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $99\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $57\%$ $39\%$ $10\%$ $17\%$ $29\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $41\%$ $77\%$ $95\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $69\%$ $50\%$ $34\%$ $18\%$ $26\%$ $36\%$
$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $81\%$ $97\%$
$(90,90)$ $100\%$ $100\%$ $92\%$ $86\%$ $99\%$ $100\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
B.2 Probit Model
Table B.6: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.150$ $0.197$ $0.063$ $0.082$ $0.042$ $0.054$
$\widehat{\beta}_{L}$ $0.055$ $0.124$ $0.013$ $0.049$ $0.007$ $0.032$
Dynamic Model
$\widehat{\beta}$ $0.159$ $0.247$ $0.067$ $0.086$ $0.043$ $0.055$
$\widehat{\beta}_{L}^{(1)}$ $0.059$ $0.134$ $0.015$ $0.052$ $0.008$ $0.032$
$\widehat{\beta}_{L}^{(2)}$ $0.062$ $0.136$ $0.016$ $0.053$ $0.008$ $0.032$
$\widehat{\rho}$ $-0.055$ $0.124$ $-0.022$ $0.054$ $-0.017$ $0.036$
$\widehat{\rho}_{L}^{(1)}$ $-0.030$ $0.105$ $-0.009$ $0.047$ $-0.007$ $0.031$
$\widehat{\rho}_{L}^{(2)}$ $-0.022$ $0.106$ $-0.003$ $0.047$ $-0.003$ $0.031$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table B.7: Empirical Size of LR, LM, and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 1
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $29\%$ $13\%$ $27\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $23\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $26\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $28\%$ $13\%$ $25\%$ $13\%$
$(60,60)$ $26\%$ $8\%$ $22\%$ $7\%$
$(90,90)$ $30\%$ $6\%$ $24\%$ $6\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.8: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 1
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $41\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $96\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $63\%$ $40\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $59\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $71\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $96\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $83\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $58\%$ $40\%$ $13\%$ $23\%$ $37\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $52\%$ $86\%$ $98\%$
$(90,90)$ $100\%$ $100\%$ $95\%$ $88\%$ $99\%$ $100\%$
Dynamic Model
$(30,30)$ $82\%$ $62\%$ $38\%$ $26\%$ $47\%$ $69\%$
$(60,60)$ $100\%$ $98\%$ $78\%$ $78\%$ $98\%$ $100\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $98\%$ $100\%$ $100\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Table B.9: Empirical Size of LM and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 2
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $69\%$ $34\%$ $59\%$ $30\%$
$(60,60)$ $74\%$ $18\%$ $62\%$ $17\%$
$(90,90)$ $71\%$ $14\%$ $60\%$ $13\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $37\%$ $29\%$ $25\%$ $26\%$
$(60,60)$ $71\%$ $18\%$ $56\%$ $16\%$
$(90,90)$ $72\%$ $14\%$ $60\%$ $13\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.10: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 2
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $97\%$ $90\%$ $77\%$ $14\%$ $13\%$ $15\%$
$(60,60)$ $100\%$ $100\%$ $96\%$ $34\%$ $70\%$ $83\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $85\%$ $98\%$ $96\%$
Dynamic Model
$(30,30)$ $95\%$ $86\%$ $74\%$ $25\%$ $28\%$ $38\%$
$(60,60)$ $100\%$ $100\%$ $93\%$ $47\%$ $84\%$ $96\%$
$(90,90)$ $100\%$ $100\%$ $100\%$ $90\%$ $99\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $96\%$ $87\%$ $73\%$ $12\%$ $13\%$ $17\%$
$(60,60)$ $100\%$ $100\%$ $96\%$ $37\%$ $69\%$ $76\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $86\%$ $91\%$ $88\%$
Dynamic Model
$(30,30)$ $94\%$ $83\%$ $70\%$ $21\%$ $27\%$ $38\%$
$(60,60)$ $100\%$ $99\%$ $92\%$ $48\%$ $81\%$ $91\%$
$(90,90)$ $100\%$ $99\%$ $99\%$ $90\%$ $94\%$ $94\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $91\%$ $83\%$ $68\%$ $10\%$ $13\%$ $19\%$
$(60,60)$ $99\%$ $99\%$ $95\%$ $39\%$ $76\%$ $92\%$
$(90,90)$ $99\%$ $99\%$ $99\%$ $88\%$ $99\%$ $99\%$
Dynamic Model
$(30,30)$ $88\%$ $75\%$ $60\%$ $21\%$ $27\%$ $39\%$
$(60,60)$ $100\%$ $99\%$ $91\%$ $49\%$ $85\%$ $98\%$
$(90,90)$ $99\%$ $99\%$ $99\%$ $91\%$ $99\%$ $100\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
C Additional Empirical Results
C.1 Sample
The data set is constructed from waves $20-30$ of PSID, covering the survey years 1987 to 1997. We selected this time span because 1997 was the last year PSID conducted annual surveys, switching to a biennial schedule in 1999. Each survey wave includes questions about employment status in the prior calendar year, such as, ‘About how many weeks did (he/she) work on that job last year?’ and ‘During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?’ We use responses to these questions to construct individual labor force participation (LFP) status for the prior year. Thus, wave 30 (1997) provides LFP data for 1996, wave 29 (1996) for 1995, and so on. In our final sample, labor force participation status is observed from 1987 to 1996, with lagged participation status observed from 1986 to 1995.
We also restrict our sample to individuals whose LFP status changed at least once during the sample period. Individuals whose labor supply status never changed during the sample period are not informative for our analysis, because the dependent variable could be perfectly predicted by setting the individual fixed effects in the intercept of our models to plus or negative infinity.
The final sample consists of a panel of $N=90$ individuals over $T=10$ years. Table C.1 presents the summary statistics. Figure C.1 shows the average percentage of individuals with given current and past LFP status. Figure C.2 plots the LFP transition rates by year.
Table C.1: Descriptive Statistics
Variable Mean Std. Dev. Min Max
Labor Force Participation 0.482 0.500 0.000 1.000
Lagged Participation 0.458 0.498 0.000 1.000
Number of Children 2.24 1.098 1.000 6.000
  • Notes: Summary statistics are computed based on a sample of 90 individuals observed from 1987 to 1996. All variables have 900 observations. Data source: PSID 1987 - 1997.
Figure C.1: Labor Force Participation Contingency Table. Each cell contains the percentage of individuals with given current and past labor force participation status.
Figure C.2: Labor Force Participation Status Transition Rates By Year
C.2 Extra Empirical Results
Table C.2: Estimated Heterogeneous Effects on Labor Force Participation
10th 25th 75th 90th
Mean Median Std. Dev. Percentile Percentile Percentile Percentile
$\alpha_{i}$ $0$0.975 6.203 -6.909 -2.443 4.292 6.774
$\gamma_{t}$ 0.000 0.146 0.773 -1.151 -0.466 0.638 0.979
$\beta_{it}$ 0.126 -0.504 6.250 -6.835 -2.628 4.050 6.982